### v1

# TeeRollup: Efficient Rollup Design Using Heterogeneous TEE

Abstract
Rollups have emerged as a promising approach to improving blockchains’ scalability by offloading transactions execution off-chain. Existing rollup solutions either leverage complex zero-knowledge proofs or optimistically assume execution correctness unless challenged. However, these solutions have practical issues such as high gas costs and significant withdrawal delays, hindering their adoption in decentralized applications. This paper introduces TeeRollup, an efficient rollup design with low gas costs and short withdrawal delays. TeeRollup employs Trusted Execution Environments (TEEs)-supported sequencers to execute transactions, requiring the blockchain to verify only the TEEs’ signatures. TeeRollup is designed under a realistic threat model in which the integrity and availability of sequencers’ TEEs may be compromised. To address these issues, we first introduce a distributed system of sequencers with heterogeneous TEEs, ensuring system security even if a minority of TEEs are compromised. Second, we propose a challenge mechanism to solve the redeemability issue caused by TEE unavailability. Furthermore, TeeRollup incorporates Data Availability Providers (DAPs) to reduce on-chain storage overhead and uses a laziness penalty game to regulate DAP behavior. We implement a prototype of TeeRollup in Golang, using the Ethereum test network, Sepolia. Our experimental results indicate that TeeRollup outperforms zero-knowledge rollups (zk-rollups), reducing on-chain verification costs by approximately 86% and withdrawal delays to a few minutes.
Index Terms:
Blockchain, Rollup, Scalability, Trusted Execution Environment
I Introduction
Since the first adoption in Bitcoin [1] in 2008, blockchain technology has experienced remarkable growth across extensive decentralized applications. However, as the number of application requests increases, blockchains have countered severe congestion due to their low throughput, i.e., scalability issues [2]. To address these challenges, rollups [3, 4], which offload a substantial portion of transaction execution off-chain, have garnered significant attention and research efforts [5]. As of April 2024, statistics from L2BEAT [6] reveal 24 rollup projects in the market, collectively holding a locked value of approximately 37 billion USD. Among these, Arbitrum [7] alone boasts 16.5 million active accounts [8].
In Rollups, the blockchain (referred to as main chain) must verify the results of off-chain transaction execution to ensure security. Current rollup solutions can be classified into two categories based on verification methods: zero-knowledge rollups (i.e., zk-rollups) [9, 10, 11], which utilize zero-knowledge proofs to validate execution, and optimistic rollups (i.e., op-rollups) [4, 7], which assume execution is valid unless challenged during a dispute period. However, these solutions come with practical issues. First, zk-rollup verification on the main chain incurs substantial gas costs, and its proof generation is time-consuming. Besides, the complex cryptographic computations of zk-rollups pose challenges for the development, deployment, and support of smart contracts. Second, the dispute period in op-rollups introduces a considerable delay for withdrawing funds. For instance, Arbitrum and Optimism have a withdrawal time of one week [7, 4].
Figure 1: An overview of TeeRollup architecture. Clients hand up transactions to the Rollup and the Rollup updates the latest state on the Ethereum.
To address these limitations, this paper presents TeeRollup, an efficient rollup design with fast withdrawal and low on-chain cost. TeeRollup leverages nodes equipped with TEEs to execute transactions off-chain, requiring the main chain to verify only the TEE-generated signatures, as shown in Fig. 1. Users can use remote attestation to verify the code within TEEs, during which a key pair is built within the TEE. The integrity of TEEs ensures correct execution, preventing malicious nodes from modifying user account status.
Unfortunately, directly using TEEs introduces challenges due to their vulnerabilities and unavailability. First, prior research has identified vulnerabilities to compromise TEEs’ integrity and confidentiality [12, 13]. To address compromised TEEs, TeeRollup adopts a distributed system of nodes with heterogeneous TEEs (e.g., Intel SGX [14], Intel TDX [15], AMD SEV [16], ARM TrustZone [17] and Hygon CSV [18], etc.). Different TEE architectures and manufacturers make it challenging to compromise a majority of TEEs. In addition, we leverage on-chain registration and attestation, avoiding the mutual attestation between heterogeneous TEEs.
Second, the I/O of TEEs is manipulated by their hosts (i.e., hypervisors or software systems), allowing malicious nodes to drop messages to and from their TEEs. Since the number of nodes in TeeRollup is small (e.g., tens of entities), there may not be enough honest nodes to ensure system availability. If TEE availability is disrupted, clients cannot redeem their funds promptly, leading to the locking of user funds on the blockchain. To mitigate this redeemability issue, TeeRollup introduces a challenge mechanism on the blockchain, enabling clients to redeem their deposits without relying on TEE-dependent operations. TeeRollup also incorporates data availability providers (DAPs) to store metadata off-chain. We introduce a collateral scheme to punish dishonest behaviors and a laziness penalty game to incentivize DAPs participation.
We build a prototype of TeeRollup with Golang [19], developing the on-chain smart contracts using Solidity 0.8.0 [20] and deploying them on the Ethereum test network, Sepolia [21]. Ethereum’s Virtual Machine (EVM) provides Turing completeness [22], fulfilling the requirements for managing the rollup. Since many TEE platforms like Intel TDX and ARM CCA currently are not yet supported by off-the-shelf CPUs, we use AMD SEV as the TEE platform [23] in our experiments. These experiments can easily be extended to other virtual machine (VM)-based TEEs.
Contributions. The contributions of this paper are listed as follows.
  • $\bullet$
    We introduce TeeRollup, an efficient rollup design that leverages TEEs for off-chain transaction execution. TeeRollup utilizes a group of heterogeneous TEEs, designed to tolerate the compromise of up to half of the TEEs. Additionally, TeeRollup incorporates data availability providers to reduce on-chain storage costs.
  • $\bullet$
    We address the redeemability issue stemming from TEEs unavailability. Our challenge mechanism ensures the redeemability of user funds, even during periods of system crash because of TEEs unavailability. Besides, we implement penalization mechanisms to ensure the diligent operation of data availability providers.
  • $\bullet$
    We conduct experimentation on the prototyped TeeRollup to evaluate its performance compared to zk-rollup and op-rollup. Our experimental results demonstrate that the transaction fee in TeeRollup is significantly lower at 0.006 USD compared to StarkNet’s 0.043 USD, showcasing the efficiency of our approach. Moreover, TeeRollup maintains a normal withdrawal time of a few minutes, consistent with StarkNet and better than op-rollup schemes such as Optimism.
Roadmap. We introduce related work in Sec. II and system model and goal in Sec. III. We present the system design in Sec. V, followed by the security analysis in Sec. VI. The system performance is evaluated in Sec. VII. Finally, the paper is concluded in Sec. VIII.
II Background and Related Work
II-A Rollup Solutions
Rollup solutions aim to alleviate the blockchains’ limitations including low throughput, high fees, and network congestion by offloading transactions processing off-chain [24]. Rollup networks are composed of sequencers who process transactions outside the main chain. However, the main chain needs to verify the correct execution of the transactions. Current rollup schemes can be divided into two categories based on their verification methods: zk-rollups and op-rollups [25]. A comparison between TeeRollup (proposed in this paper) and prior works is provided in Table I.
Zk-rollup models, as embodied by implementations like StarkEx [26], generate the zero-knowledge proofs [27] for the correct off-chain execution of transactions. Moreover, the emergence of solutions like StarkNet [10] and Scroll [28], involves EVM-compatible zero-knowledge proof. This unique capability extends to the verification of contract (not just transactions) execution correctness, encompassing both input and output validity in the process. Thomas et al. further introduce an on-demand zk-rollup creation service [29], which allows several zk-rollups to co-exist as groups on the same smart contracts, and application examples for Internet of Everything (IoE) [30]. However, the on-chain verification of zero-knowledge proofs in zk-rollup systems necessitates a notable gas expenditure.
Op-rollup solutions, implemented by Optimism [4] and Arbitrum [7], pivot towards an optimistic approach for transaction execution. These solutions optimistically assume that transactions are executed correctly by sequencers unless a challenger disputes the execution results. The dispute can be solved on the main chain. The advantage here lies in omitting on-chain verification, reducing the gas cost. Nevertheless, op-rollups significantly extend the withdrawal time, as they require sufficient time for challengers to verify the proof, which can take up to one week.
II-B Rollup Data Availability
One critical aspect of rollup solutions is ensuring data availability, which refers to the accessibility and integrity of transaction data necessary for maintaining the security and trustworthiness of the system [31]. Storing complete data on the main chain inherits the security and data integrity of the main chain, but faces challenges such as network congestion and high gas costs. For instance, storing a 256-bit integer on the Ethereum smart contract STORAGE field costs 8K gas (0.46 USD) [20]. Therefore, it is unpractical to simply permanently store the complete data of the rollup on the main chain. StarkEx [26] and StarkNet [10] address data availability issues by including aggregated compressed transactions in the CALLDATA field of the main chain, which reduces the gas cost [32]. However, this method does not allow for permanent storage of transactions and consumes more gas compared to only recording the state root on the main chain.
II-C Trusted Execution Environment
Trusted Execution Environments (TEEs) aim to run applications in a secure environment without leaking secrets to an adversary who controls the computing infrastructures. Specifically, TEEs provide an isolated area (known as enclave) to run code and the attestation mechanism to prove the correctness of computation. Influential TEE implementations include Intel SGX [14], Intel TDX [15], ARM TrustZone [17], ARM CCA [33], AMD SEV [23] and Hygon CSV [18].
TABLE I: Rollup Solution. The Dec, EVM Comp, and Data Avail are short for Decentralized, EVM Compatibility, and Data Availability, respectively. Decentralized property means that the transaction cannot confirmed by a single sequencer. The EVM compatibility denotes the ability to support EVM. For Data Availability, Main Chain and Delegated means the data availability are provided by the main chain and a third party, respectively. Withdrawal time signifies the waiting time for users to retrieve the funds they locked in the main chain.
Rollup Solution Dec EVM Comp Data Avail Withdrawal time
StarkEx [26] Main Chain Few minutes
StarkNet [10] Main Chain Few minutes
Scroll [28] Main Chain Few minutes
Optimism [4] Main Chain One week
Arbitrum [7] Main Chain One week
TeeRollup Delegated Few minutes
Recently, TEEs have been widely used in blockchain designs to enhance security, privacy, and performance [34, 35, 36, 37, 38, 39, 40]. Teechain [35] establishes a payment system under TEE protection, while Bool Network [41] employs TEE to ensure the privacy of secret key components within cross-chain platforms. Tommaso et al. [36] and Xu et al. [37] leverage the TEE for the correctness and privacy of off-chain execution, respectively. These works assume that TEEs provide integrity and confidentiality guarantees but do not ensure the availability of the running service. Specifically, an adversary with full control of the operating system (OS) can arbitrarily launch, suspend, resume, terminate, or crash a TEE at any time. They can also delay, replay, drop, and inspect the messages sent to and from TEEs. Furthermore, the adversary cannot know the private key $sk_{i}$ inside the TEE. However, the software inside TEEs is off-limit.
However, many studies indicate that TEEs are vulnerable to various attacks such as side-channel attacks [42], unprotected I/O [43], and ASID abuses [44]. Thus, in our work, we assume TEEs can be compromised, i.e., no integrity and confidentiality properties, which differs from prior work with the perfect assumption of TEEs [35, 36, 37, 38, 39, 40]. Particularly, the adversary can steal the private keys for signing messages from the compromised TEE. To tolerate malicious behaviors of compromised TEEs, one approach is using heterogeneous TEEs to jointly do computation. This heterogeneity can arise from using TEEs from different vendors, different types of processors, or different levels of security requirements [45]. Since different TEE platforms adopt different hardware and software architectures, it is difficult to breach their security simultaneously. Besides, when a TEE is compromised, manufacturers will timely resolve these issues, making security compromise of all TEEs nearly impossible [46, 47].
III Problem Statement and System Model
In this section, we first present the problem statement and the system model. Then, we introduce the associated challenges and necessary preliminaries. Table II lists the frequently used notations in this paper.
III-A Problem Statement
Consider a rollup system of $n$ sequencers (i.e., nodes with TEEs). The rollup system issues tokens (i.e., TTokens), and provides deposit, transfer, and redeem functions for clients. A client on the main chain $\mathcal{M}$ with account $c_{addr}$ can enter rollup by depositing on the account $T_{addr}$ held by the rollup. Then, the rollup issues the TTokens and locks the deposit on the account $T_{addr}$. Clients can submit transactions to sequencers to transfer inside the rollup. Sequencers handle transactions and update the state of the rollup to the blockchain $\mathcal{M}$. The processing of transactions inside rollup can be defined as $state_{h+1}\leftarrow execute(state_{h},txs)$, where $txs$ is a batch of transactions.
TABLE II: Summary of notations.
Term Description Term Description
$\mathcal{M}$ main chain $state_{h}$ state root of TeeRollup
TSC TeeRollup smart contract $h$ height of $state_{h}$
MSC Manager smart contract $n$ Number of sequencers
TToken Tokens of TeeRollup $p_{i}$ Sequencers in TeeRollup
$T_{addr}$ TSC account on $\mathcal{M}$ $\eta_{i}$ Enclave for sequencer $p_{i}$
$\tau_{c}$ Timer for resolve challenge $(pk_{i},sk_{i})$ Key pair for enclave $i$
III-B System Model
We consider a system of $n$ sequencers, denoted by the set $\mathcal{G}=\{p_{1},p_{2},...,p_{n}\}$. We follow the existing model [48], where each sequencer $p_{i}$ is equipped with one type of TEE platform (e.g., Intel SGX, Intel TDX, etc), denoted by $\eta_{i}$. We assume at most $f$ sequencers’ TEEs are compromised at any time, while $n-f$ sequencers’ TEEs are uncompromised. Here, since TEE machines adopt different architectures and are produced by different manufacturers, it is difficult to breach the security of $f$ ones at the same time. The parameter $f$ can be adjusted. There is a public/private key pair of the sequencer $p_{i}$, denoted by $(pk_{i},sk_{i})$, in which the private key is only stored and used inside TEE.
Threat model. We model the malicious sequencers as Byzantine adversaries, i.e., they can behave in arbitrary ways. The threat model of sequencers consists of three cases: the honest sequencers with/without compromised TEEs, the malicious sequencers with compromised TEEs, and the malicious sequencers with uncompromised TEEs.
  • $\bullet$
    Honest sequencers with/without compromised TEE. We assume the honest sequencers with their TEEs do not deviate from the protocol.
  • $\bullet$
    Malicious sequencers with uncompromised TEEs. We assume the malicious sequencers have full control over the operating system of their TEEs, including root access and control over the network. The sequencers can arbitrarily launch, suspend, resume, terminate, and crash TEEs at any time. Besides, the sequencers can delay, replay, drop, and inspect the messages sent to and from TEEs, i.e., manipulating input/output messages of TEEs. In other words, the TEE cannot guarantee availability due to these manipulations.
  • $\bullet$
    Malicious sequencers with compromised TEEs. We assume the sequencers have full control of the enclave and know the private key $sk_{i}$ inside the compromised TEE. In this case, there are no such integrity and confidentiality guarantees for TEE. We assume static corruption by the adversary, where a fixed fraction of all sequencers’ TEEs is compromised.
We consider the worst case, in which all the sequencers can be malicious and controlled by a single adversary $\mathcal{A}$. Note that the rollback attacks are orthogonal to this work and can be addressed by work [49, 50, 51].
Main chain model. We assume the TeeRollup is built on the main chain $\mathcal{M}$. The main chain provides finality (i.e., once a transaction is included in a block it is considered final) [52] and enables smart contracts [53]. Smart contracts can access the current time using the method $block.timestamp$ and provide cryptographic schemes including hash computing and ECDSA encryption [20].
III-C System Goals
In this work, we aim to design a rollup to execute transactions and record the state of the rollup on the main chain, which satisfies the following properties:
  • $\bullet$
    Trust-minimized participants: The client does not need to trust the other party or an intermediary to ensure that the transaction is executed honestly.
  • $\bullet$
    Redeemability: Any client can redeem TTokens in TeeRollup without requiring a third party (i.e., redeemability relies only on the secure operation of the main chain).
  • $\bullet$
    Liveness: Any transactions sent to TeeRollup are either executed or settled on the main chain.
III-D Design Challenges
To better outline the TeeRollup design, we first introduce a strawman solution (referred to as SRollup) that leverages TEE to shield the sequencer from malicious behavior. In SRollup, a single sequencer collects and executes the transactions from clients and submits the execution result to the main chain. While SRollup always provides the fundamental functions of issuing, transferring, and redeeming TTokens, it does not achieve all the goals defined in III-C. This is because the SRollup does not consider the compromised TEE and potential malice of sequencers and DAPs.
Compromised TEE. As mentioned in  Sec. III-B, the TEE can be compromised and sequencers can acquire the secret key of the compromised TEE. Since the state update in the system is confirmed on the main chain, TEE generates a new state with the latest state on the chain as the previous state. Compromised TEE may forge false transactions or incorrectly execute transactions to generate a new state. Therefore, a single TEE cannot be directly trusted by the system. To counteract the compromise of a certain proportion of TEEs, TeeRollup employed a group of sequencers equipped with independent and heterogeneous TEEs. The sequencer checks the validity of the state and signs it inside the TEE. Then, if the sequencer collects $f+1$ types of TEEs’ signatures, it can submit the state with signatures to the main chain. Besides, TeeRollup adopts the on-chain attestation when registration, avoiding the attestation of heterogeneous TEEs.
Malicious sequencers. According to our threat model (see Sec. III-B), the operating system for TEE is fully controlled by the sequencer. Malicious sequencers can crash the TEE, or censor clients’ transactions and filter out some of them. At this point, if the TEEs are unavailable, TeeRollup cannot process any clients’ requests and clients are unable to redeem from TeeRollup, thereby destroying its Liveness and Redeemability. Therefore, TeeRollup presents a challenge mechanism for clients to solve the redeemability issue, and make an on-chain settlement when TEE does not provide service. Once the client thinks that his transaction has been reviewed, that is, it has not been processed for a long time, he can submit the transaction data to the TSC through the challenge method provided on the TSC. Once a challenge is initiated in TSC, the timer is also set. If the sequencer does not respond to the challenge in time, the system is considered unavailable and enters the on-chain for settlement. Therefore, the verifier chooses to respond to the challenge in time to prevent the whole system from settlement. Moreover, this also solves the problem of redeeming client funds when the system is not available.
Lazy data availability providers. The states of TeeRollup are recorded in TSC, and its integrity is guaranteed by the main chain (see Sec. III). Therefore, the correctness of the metadata provided by DAPs can be verified by the $state$ recorded in TSC (i.e., DAPs cannot forge account balance, transaction data, or other metadata). However, the DAPs can show laziness, discarding the data they are supposed to store. Thus, clients and sequencers cannot get metadata of the $state$ on TSC. TeeRollup introduces collateral and laziness penalty game to incentivize DAPs to behave diligently. DAPs need to provide the collateral when entering the network and once a DAP acts dishonestly, some or all of its collateral will be deducted.
IV Preliminaries
[…]
Merkle Tree. Merkle tree is a binary tree data structure in which each leaf node contains the hash of a data block (e.g., the key-value pair), and each non-leaf node contains the hash of its child nodes. Constructed bottom-up, they generate a single hash at the root, representing the entire dataset’s integrity. The root of a Merkle Tree is integrated into each block, serving as a comprehensive state digest, which enables sequencers to maintain and verify the state digest efficiently. The Merkle Tree allows leaves to generate a Merkle Path for verifying specific key-value pairs under a root, denoted as $\delta\leftarrow root.proof(\left<k,v\right>)$. Anyone can verify that this pair of key-value through $verifyproof(\delta,\left<k,v\right>)$.
Figure 2: An overview of TeeRollup architecture.
Trusted Execution Environments (TEEs). We follow prior work [54] to model the functionality of attestable TEEs. Each TEE is initialized with a key pair $(sk,pk)\leftarrow\Sigma.keyGen(1^{\lambda})$ generated by its manufacturer, where $\Sigma$ is the signature algorithm. Here, $sk$ represents the TEE’s internal master secret key, while $pk$ is the corresponding public key. The ideal functionality of a TEE offers the following Application Interfaces (APIs) for a program code $prog$:
  • $\bullet$
    $pk\leftarrow getpk()$: Retrieve the TEE’s public key $(pk)$.
  • $\bullet$
    $eid\leftarrow install(prog)$: Install the program $prog$ as a new enclave within the TEE, assigning it a unique enclave ID, $eid$.
  • $\bullet$
    $(outp,\sigma)\leftarrow resume(eid,inp)$: Resume the enclave with ID $eid$ to execute program $prog$ using the input $inp$. $\sigma$ is the TEE’s endorsement, confirming that $outp$ is the valid output.
For uncompromised TEE, with the acquired $m_{pk}$, the verification of the output $outp$ is trusted by users when $\Sigma.verify(\sigma,eid,prog,outp)=1$, indicating the successful verification of the signature. Besides, we assume an attestation API for TEE to generate an attestation quote $\rho\leftarrow attest(eid,prog)$ proving that the program $prog$ has been installed in the enclave $eid$. And $\rho$ can be verified through method $verifyquote(\rho)$.
V TeeRollup Design
V-A Overview
Fig. 2 shows the architecture of TeeRollup, a rollup based on TEE. TeeRollup considers four roles: the TeeRollup Smart Contract (TSC) responsible for record-keeping on the main chain, the Manager Smart Contract (MSC) for managing sequencers, the TEE-equipped sequencer committee processing off-chain transactions, and the data availability provider (DAP) storing metadata for TeeRollup, as shown in Fig. 3.
  • $\bullet$
    TeeRollup smart contract (TSC). TSC records the TeeRollup state proof in the form of a state root (rather than complete metadata) to reduce on-chain storage cost.
  • $\bullet$
    Manager smart contract (MSC). MSC manages registered TEEs of sequencers, especially assigning an ID for them and recording the private key $m_{pk}$ inside TEE.
  • $\bullet$
    Sequencer. Sequencer primarily collects and executes transactions from clients, submits the state root to TSC, and sends metadata to DAP.
  • $\bullet$
    Data availability provider (DAP). DAP accepts and stores the metadata sent by sequencers to ensure access to system metadata at any time. The metadata includes the state of the system (i.e., account balances), the update history, and transaction data.
There are four key procedures in TeeRollup. First, during the initialization phase, the service provider deploys both the TSC and the MSC on the blockchain $\mathcal{M}$. Subsequently, sequencers can register their TEEs on the MSC (see Sec. V-C). Second, clients deposit funds on the TSC and request the issuance of TTokens in TeeRollup, which can be used for transfers. Sequencers receive transactions from clients, process these transactions, and generate an updated state root. Then, the sequencers broadcast the new state root along with metadata and collect votes. Once enough votes are collected, the sequencers submit the state root of TeeRollup to the TSC and send the metadata to the DAPs (see Sec. V-D). Third, if the transactions of clients are not processed in time, clients can challenge the TSC to enforce the transaction. Even if the whole system cannot make any response, they can also get a refund directly on the TSC to avoid the loss of property (see Sec. V-E). Fourth, to better regulate the behavior of DAPs and prevent laziness, we use the incentive method of collateral plus laziness punishment (see Sec. V-F).
V-B Data Structure
State root format. In TeeRollup, the state root acts as an abstract of the metadata and anyone can verify the transactions, balance, and other information of TeeRollup using this abstract. The state root $state_{h}$ on height $h$ has the following structure:
$state_{h}:=\left<h,H(state_{h-1}),AR_{h},H(txs_{h})\right>$
where $H(state_{h-1})$ refers to the hash of the previous state root $state_{h-1}$, $AR_{h}$ is the Merkle root of the Account Tree of TeeRollup, and $H(txs_{h})$ is the hash of the transactions list executed to generate the $state_{h}$. Account Tree is a Merkle Tree that stores the clients’ account in TeeRollup, denoted as $A_{h}$. Account Tree $A_{h}$ consists of the account address and the balance for all accounts. Note that clients of TeeRollup use the same account address and private key with the main chain.
In TeeRollup, the state roots utilize the chain structure, which has been adopted by Bitcoin [1], Ethereum [22], and some BFT protocols [55]. Every state root contains the hash of the previous state root, and the state root can be indexed by height (i.e., the distance from the initial state root). In TeeRollup, there is only one state root at each height. Once the latest state root is obtained, the correctness of any historical state can also be verified.
Vote and certificate. To achieve collaboration among multiple sequencers, TeeRollup adopts a competition for the right to submit states. Specifically, any sequencer can become the leader, broadcast the state root, and collect votes. However, the TSC will only accept the first state that arrives at one height. A vote $v_{i}$ from sequencer $p_{i}$ of state root $state_{h}$ has the following structure:
$v_{i}:=\left<H(state_{h}),pk_{i},\sigma_{i}\right>$
where $H(state_{h})$ is the hash of $state_{h}$, $pk_{i}$ is the public key for sequencer $p_{i}$, and $\sigma_{i}$ is a signature created by the sequencer $p_{i}$ over $H(state_{h})$. Here, the signature is generated inside the TEE of sequencer $p_{i}$ with the private key $sk_{i}$. If there is a set of $f+1$ type of TEEs’ signature, forming the quorum certificate (QC) for the state root. Here, a QC can be implemented as multi-signatures or aggregated signatures. If a state root attaches the QC, it is acceptable for the TSC. Every sequencer keeps updating the local state root in line with TSC. In TeeRollup, a new state root $state_{h+1}$ that meets the following conditions can be certificated by the TSC. First, the height of the lasted state root in TSC is $h$. Second, $state_{h+1}$ contains the hash value of $state_{h}$, which corresponds to the lasted state root in TSC. Third, the QC for $state_{h+1}$ is valid.
Figure 3: High overview of Issue, Deal and Redeem process.
V-C Sequencers Registration and Configuration
TeeRollup introduces the MSC to manage the sequencer committee, which provides the registration and configuration for sequencers.
Enclave registration. Sequencers equipped with TEEs can contribute to TeeRollup by creating an enclave $\eta_{i}$ and registering it on the MSC. Here, we present an exemplary representation of a sequencer $p_{i}$ registering on MSC for clarity. Sequencer $p_{i}$ creates an enclave $\eta_{i}$, and installs the TeeRollup program $prog$ inside $\eta_{i}$, which is mandated to run correctly within $\eta_{i}$. Upon the creation of $\eta_{i}$, an asymmetric key pair $(pk_{i},sk_{i})$ is generated. Note that the public key is also an account address for the main chain [22]. The secret key $sk_{i}$ is securely stored within $\eta_{i}$, accessible only by the TeeRollup program running inside $\eta_{i}$, while the public key $pk_{i}$ is returned as output to the sequencer $p_{i}$. Subsequently, $\eta_{i}$ generates a proof $\rho_{i}$ asserting its execution of the TeeRollup program and control over the $sk_{i}$ corresponding to $pk_{i}$. Then, sequencer register $\eta_{i}$ by invoking Register and forwarding $\eta_{i},\rho_{i},pk_{i}\right>$ to the TSC. The TSC verifies that $verifyquote(\rho_{i})=1$ [54, 56]. Upon successful verification, TSC adds the sequencer $p_{i}$ with $pk_{i}$ to the sequencer list. The registration process guarantees the execution of the $prog$ for all registered enclaves and the confidentiality of the secret key $sk_{i}$. Thus, the heterogeneous TEE in the system can trust the attestation of other TEEs on TSC, and this key pair $pk_{i},sk_{i}\right>$ can be used for generating signatures between TEEs. There is no necessity to re-attest enclaves in subsequent protocol steps.
Sequencer configuration. MSC provides the management for sequencers, and sequencers can register in or exit the TeeRollup freely. MSC maintains a sequencers list, storing their public key and the committee strategy of the TSC. In this paper, we adopt a $f+1$ voting strategy (i.e., if the multi-signature contains signatures from more than $f+1$ type of TEEs of sequencers in the committee, it is considered valid).
V-D Normal-Case Operations
As shown in Fig. 3, TeeRollup provides issue, transfer, and redeem functions for clients.
V-D1 Issue
Client deposits in the TSC and the TeeRollup issues the corresponding amount of TTokens for the client. Note that the exchange rate between the currency of blockchain $\mathcal{M}$ and TToken is beyond the scope of this paper, which is set to be 1.
Step 1. The client invokes the Deposit on the TSC, denoted as a tuple of $s_{addr},v\right>_{\sigma}$, where $s_{addr}$ is the client’s address and $v$ is the deposit value (greater than zero), and $\sigma$ is the signature of the client.
Step 2. The enclave $\eta_{i}$ within the sequencer $p_{i}$ monitors the $deposit$ transaction on TSC. Enclave $\eta_{i}$ issues TTokens and changes the balance of account $s_{addr}$. Note that the client holds the same address in TeeRollup as in the blockchain $\mathcal{M}$. Enclave $\eta_{i}$ batches multiple transactions together to optimize efficiency and executes these transactions, generating the new state root $state_{h+1}$ and a $lock$ transaction denoted as $id\right>$, where $id$ is the number of the $deposit$. The sequencer $p_{i}$ sends the $state_{h+1}$ and $lock$ with the metadata to other sequencers, who vote for the $state_{h+1}$.
Step 3. After collecting $f+1$ types of TEEs’ votes, the sequencer $p_{i}$ generates the QC for $state_{h+1}$ and sends $state_{h+1}$ and the QC to the TSC. TSC verifies the validity of $state_{h+1}$ and QC (detailed in V-B). Note that TSC only accept the first validate $state_{h+1}$ in height $h+1$. If successful, TSC records $state_{h+1}$ on the main chain. Besides, TSC locks the $deposit$ and the client cannot withdraw the $deposit$ unless exit the TeeRollup. Note that the update of $state_{h+1}$ and the lock of $deposit$ are invoked in one call, such that they can be confirmed together.
step 4. After $state_{h+1}$ and the lock transaction is confirmed on the main chain, the sequencer $p_{i}$ returns the execution result to the clients and sends the metadata to the DAPs.
V-D2 Transfer
Client transfers their TTokens by sending transactions to the sequencer.
Step 1. A client submits a transaction $tx$ to the sequencers. $tx$ can be denoted by tuple of $s_{addr},r_{addr},v\right>_{\sigma}$, where $s_{addr}$ is the sender’s address, $r_{addr}$ is the receiver’s address, $v$ is the transferred value (greater than zero), and $\sigma$ is the signature of the sender.
Step 2. The execution of $tx$ changes the balance of account $s_{addr}$ and $r_{addr}$. Enclave $\eta_{i}$ batches $tx$ with other transactions, executes them, and generates the new state $state_{h+1}$ signed with its private key $sk_{i}$. Then, sequencer $p_{i}$ broadcast $state_{h+1}$ with $state_{h}$ and transactions $txs$ to others and waiting for their votes.
Step 3. After collecting $f+1$ types of TEEs’ votes, the sequencer $p_{i}$ generates the QC for $state_{h+1}$ and sends $state_{h+1}$ and the QC to the TSC. TSC verifies that QC is valid, and then records $state_{h+1}$.
Step 4. If $state_{h+1}$ is confirmed on the main chain, the sequencer $p_{i}$ returns the execution result to the clients and sends the metadata to the DAPs.
V-D3 Redeem
Client burn TTokens in TeeRollup and TSC refund to the Client. Clients burn TTokens by transferring TTokens to the Account $A_{0}$, which is the burning account. Account $A_{0}$ can only receive tokens, but cannot transfer to others, and tokens transferred to the $A_{0}$ are considered burned.
Step 1. Clients submit requests to the sequencer $p_{i}$, and the request can be denoted as $s_{addr},A_{0},v\right>_{\sigma}$, where $s_{addr}$ is the client’s address, $v$ is the transferred value (greater than zero), and $\sigma$ is the signature of the client.
Step 2. Enclave $\eta_{i}$ batches $tx$ with other transactions, executes them, and generates the new state $state_{h+1}$ with a $refund$ transaction for the TSC, denoted as $s_{addr},v\right>$. Enclave $\eta_{i}$ signs $state_{h+1}$ and $refund$ with its private key $sk_{i}$. Then, the sequencer $p_{i}$ sends $state_{h+1}$ with $refund$ to others and waits for their votes.
Step 3. After collecting $f+1$ types of TEEs’ votes, the sequencer $p_{i}$ generates the QC for $state_{h+1}$ and sends $state_{h+1}$ and the QC to the TSC. If QC is verified to be valid, TSC records $state_{h+1}$ on the main chain and refunds the currency with the value of $v$ to $s_{addr}$.
Step 4 If $state_{h+1}$ and $refund$ transaction are confirmed on the main chain, sequencer $p_{i}$ returns the execution result to the clients and sends the metadata to the DAPs.
1:Function Init
2:    ContractState $\leftarrow$ Active
3:    $Chal\leftarrow\emptyset$
4:    $Dep\leftarrow\emptyset$
5:
6:Function Deposit ($M_{addr},T_{addr},v$)
7:    Require ContractState = Active
8:               $\wedge$ $value>0$
[…]
10:    $Dep_{id}.sender\leftarrow M_{addr}$
11:    $Dep_{id}.value\leftarrow v$
12:    $Dep_{id}.solved\leftarrow false$
13:    Trigger Deposit event
14:
15:Function UpdateState($QC,S$)
16:    Require ContractState = Active
17:               $\wedge$ $S.h=height$
18:               $\wedge~{}S.preHash=H(State_{h}eight)$
19:               $\wedge$ VerifyMutiSig($QC$)
20:    $height\leftarrow t+1$
21:    $State_{height}.root\leftarrow S$
22:
23:Function StartChallenge ($M_{addr},tx,pledge$)
24:    Require ContractState = Active
25:               $\wedge$ $pledge\geq Pledge_{C}$
26:    $id\leftarrow$ H($M_{addr},tx,block.timestamp$)
27:    $Chal_{id}.tx\leftarrow tx$
28:    $Chal_{id}.startChal\leftarrow block.timestamp$
29:    Trigger Challenge event
30:
31:Function ResloveChallenge ($id,QC$)
32:    Require ContractState = Active
33:               $\wedge$ VerifyMutiSig($QC$)
34:    Delete($Chal_{id}$)
35:
36:Function SettleRollup ($id$)
37:    Require ContractState = Active
38:               $\wedge$ $block.time-Chal_{id}.startChal>T_{w}$
39:    ContractState $\leftarrow$ Rrozen
40:    Trigger Settle event
41:
42:Function SettleWithdraw ($T_{addr},M_{addr},b,\delta,\sigma$)
43:    Require ContractState = Frozen
44:               $\wedge$ Verify($T_{sddr},\sigma$)$=1$
45:               $\wedge$ VerifyProof($\delta,b,T_{sddr}$)$=1$
46:    Refund($M_{addr},b$)
Algorithm 1 Challenge Algorithm of TRollup Smart Contract
V-E Challenge Mechanism
To address the liveness and redemption issue, we propose a challenge mechanism deployed in TSC. In TeeRollup, clients can initiate a challenge in TSC if their request is not processed. If there is no response from the sequencers, clients can also withdraw their deposit from TSC with the proof of balance provided by DAP.
Specifically, clients can invoke the StartChallenge to initiate a challenge on TSC. TSC starts a $challenge$ and gives it an $id$ (calculated by the hash of the sender address, transaction, and block timestamp). Then, TSC sets a timer with the current block timestamp as the challenge start time, and triggers an $challenge$ event with $s_{addr},tx,block.timestamp\right>$. If the enclave receives the $challenge$ from the sequencer, who monitors the challenge event, the enclave will execute the $tx$ in the next batch and vote to respond to it before $\tau_{w}$ passes. Therefore, when the sequencers provide the multi-signature from the enclaves for the $tx$, it proves that the enclaves have received the transaction and executed it. Note that, to provide the sequencers enough time to process the transaction and reply, $\tau_{w}$ is set long enough for a few days or a week.
On the other hand, if the sequencer drops the request, and cannot generate the response to the challenge in time, clients can invoke the SettleRollup($id$) to settle the TSC. TSC verifies whether the current block time has exceeded the start time of the challenge plus the waiting time $\tau_{w}$. If the verification is passed, the contract state of TSC becomes frozen. At this time, sequencers can no longer update the state root, and deposit is also not allowed, while the client can withdraw their balance. Note that, to prevent the client’s malicious challenge (for example, they do not submit a transaction $tx$ to the sequencer, but carry out a challenge on the chain), the client needs to lock collateral on TSC when initiating a challenge. If the challenge fails, the collateral will be deducted. Besides, the client also needs to submit the $s_{addr},r_{addr},v\right>$ to the TSC, so that sequencers can get the complete message of $tx$. The waiting time $\tau_{w}$ is set long enough for the sequencers to monitor the challenge from the TSC and reply.
To withdraw the balance $b$ in account $a$, the client needs to provide two proofs: (i) the client has the key of account a; and (ii) the balance in account a is b. First, according to Sec. V-D, the account address $a$ is the public key and the client can generate the signature with the private key, proving their control of account $a$. Second, the client can get the proof of balance from DAP, i.e., the Merkle proof for the balance of the client’s account. The state root recorded in TSC contains the $AccountRoot$, which is the Merkle root of the account balance. Thus, the DAPs can provide the balance $b$ and proof $\delta\leftarrow state_{h}.AccountRoot.proof(a,b)$, where $a$ is the address for the account $b$ in TeeRollup. Client generates the signature $\sigma$ for $\delta$ with the secret key of account $b$ and invoke SettleWithdraw($T_{addr},M_{addr},b,\delta,\sigma$). Finally, TSC verifies the proof of balance and the signature $\delta$. If successful, TSC refund the balance of account $b$ to the client. Thus, every client can redeem their deposit on TSC.
The above design solves the problem that the client’s transfer and withdraw (i.e., a transfer to a specific account) transactions are not executed, but the client’s deposit is also at risk of not being processed. Thus, in our design, every $deposit$ should be confirmed by the sequencers. When clients raise a TTokens request and lock the $deposit$ on the TSC, the $deposit$ is unsolved, and a timer with a countdown of $\tau_{w}$ is set. Then, sequencers process the $deposit$, issue TTokens for the client, and submit the new state root to TSC. Upon the new state root is confirmed on the main chain, sequencers respond to the $deposit$, and set it to be solved. If the $deposit$ is unsolved until $\tau_{w}$ passes, TSC can refund $deposit$ to the clients.
V-F Data Availability Provider
To minimize on-chain storage expenses, we have relocated the support for data availability off-chain (i.e., data availability providers). For a state root $state_{h}$, DAPs store the $state_{h}$, transaction lists $txs_{h}$, and Account Tree $A_{h}$, which we called the metadata of TeeRollup. Alongside the submission of updated states to the TSC, sequencers simultaneously upload the metadata to the DAPs. However, DAPs can act lazy (e.g., withholding data or lazy validation) and pretend as if metadata was stored. Therefore, to motivate DAPs to store metadata diligently, we have designed a laziness penalty game to punish the lazy DAP.
Registration. To ensure that DAPs store the metadata of the state recorded on TSC, we require DAPs to register and lock a specified collateral amount. Here, we present an exemplary representation of a DAP registering on TSC for clarity. Initially, the DAP registers by submitting $D_{addr},v\right>_{\sigma}$ to TSC, where $D_{addr}$ represents the DAP’s address (i.e., the public key), and $v$ is the collateral value. Upon receipt, TSC authenticates the signature $\sigma$ and validates whether $v$ exceeds the minimum requirement value for collateral, denoted as $C$. If a DAP successfully passes the verification, it is authorized to store the metadata of TeeRollup.
Laziness penalty game. We introduce the collateral $C$ and punishment for lazy DAPs to incentivize DAPs. We refer to this component of TSC as the laziness penalty game. Specifically, the client can initiate random data requests to the DAPs, and the node that fails to post a response loses part or all of its collateral (i.e., a slashing of collateral). Once the data request is triggered, the DAPs should respond to the TSC within time $l$, otherwise, the DAPs will suffer a slashing of collateral. The response consists of the metadata and the Merkle path proof for the metadata.
For the slashing mechanism, we follow the design in [57], which introduced an optimal slashing function for DAPs. In our design, for a system with $m$ DAPs, each DAP stores a copy of the metadata. Thus, any DAP can respond to the request and provide the complete metadata. Besides, the response from DAP can be verified by the TSC, and the invalid response is considered to be a no-response. For a specific data request, the sets $X=\left<x_{1},x_{2},...,x_{m}\right>$ represent the reply of DAP. Let $x_{i}=1$, if DAP $q_{i}$ sends a valid data response to TSC, and $x_{i}=0$ otherwise. Since this design is a special case for [57], the slashing function for $q_{i}$ is defined as follows.
$f_{i}(x)=\begin{cases}0&,i=1\\ -C&,\sum_{j=1}^{m}x_{j}<1~{}and~{}x_{i}=0\\ -W-\epsilon&,\sum_{j=1}^{m}x_{j}\geq 1~{}and~{}x_{i}=0\\ \end{cases}$
The TSC slashes the collateral $C$ put up by each DAP that has not sent a valid response before the timeout if there is no valid response. Otherwise, if there is more than one response in the TSC before the timeout, it punishes the non-responsive DAPs by a modest amount, namely $W+\epsilon$, where $W$ is the cost for the DAP to construct and send a response to TSC. Note that $\epsilon$ is a small number to prevent the DAP from responding negatively to save the cost $W$.
VI Security Analysis
In this section, we analyze the security of TeeRollup, and prove that TeeRollup satisfies the Redeemability, and Liveness properties defined in Sec. III-C.
Lemma 1.
If a malicious sequencer forges an invalid current state root $state_{h}^{\prime}\neq state_{h}$ to its protected enclave, the new state root $state_{h+1}^{\prime}$ generated by the enclave will not be accepted by TSC.
Proof.
According to the Sec. V-D, the honest enclave generates the $state_{h+1}^{\prime}$ containing the hash of $state_{h}^{\prime}$, i.e., $state_{h+1}^{\prime}.preHash\leftarrow H(state_{h}^{\prime})$. Thus, if the sequencer submits the $state_{h+1}^{\prime}$ signed by the enclave, TSC will check whether $state_{h+1}^{\prime}.preHash$ is equal to $H(state_{h})$. Obviously, since $state_{h}^{\prime}\neq state_{h}$, TSC will not accept the $state_{h+1}^{\prime}$. ∎
Lemma 2.
If a malicious sequencer with the compromised TEE forges an invalid state root $state_{h+1}^{\prime}$ with transaction list $txs_{h+1}^{\prime}$ (i.e., $state_{h+1}^{\prime}$ cannot be calculated by executing transactions $txs_{h+1}^{\prime}$ with the initial state $state_{h}$), TSC will accept the state root.
Proof.
By Lemma 1, the sequencer with the protected TEE cannot make the enclave output a forged state by providing the forged input. If the sequencer provides the correct input $state_{h}$, the enclave will only generate $state_{h+1}$ by executing $txs_{h+1}^{\prime}$. This is because the program in the enclave is protected from execution. Thus, the malicious sequencer with protected TEE cannot generate the forged state root $state_{h+1}^{\prime}$.
According to Sec. III, TeeRollup have $n$ sequencers, and up to $f$ types of TEE can be compromised. Thus, the $state_{h+1}^{\prime}$ can only be signed by at most $f$ types of compromised enclaves of sequencers, i.e., it will not be accepted by TSC. ∎
Theorem 1 (Trust-minimization).
Malicious sequencers cannot upload an invalid state accepted by the TSC.
Proof.
According to Lemma 1 and Lemma 2, malicious sequencers with uncompromised TEE or compromised TEE cannot forge invalid state roots accepted by the TSC. ∎
Theorem 2 (Redeemability and Liveness).
If a client submits a transaction $tx$ to sequencers, the transaction will finally be executed or the TSC will be settled (i.e., the deposit of all clients can be redeemed in the main chain).
Proof.
According to the challenge mechanism, the client can initiate a challenge, if its transaction is not executed for a long time. Once a challenge is initiated, there are two cases as follows.
  • $\bullet$
    Case 1. Sequencers resolve the challenge before the waiting time $\tau_{w}$ passes. In this case, at least $f+1$ enclaves of sequencers have received the transaction $tx$. Thus, at least one protected enclave will batch the transaction and execute it to generate a new state root.
  • $\bullet$
    Case 2. Sequencers fail to resolve the challenge before the waiting time $\tau_{w}$ passes. As a result, the client can invoke the Settle($id$) to settle the TSC. TSC is frozen immediately, and any client can redeem their deposit on TSC.
VII Performance Evaluation
We evaluate TeeRollup in terms of its on-chain verification costs, throughput, and latency of processing off-chain transactions. We consider two state-of-the-art counterparts, zk-rollups [3] and op-rollups [4]. With various experiments, we answer the following questions:
  • $\bullet$
    Q1: How does TeeRollup perform in the on-chain cost?
  • $\bullet$
    Q2: How does TeeRollup perform in throughput and latency?
  • $\bullet$
    Q3: How does TeeRollup perform compared to zk-rollups and op-rollups?
TABLE III: System Parameters.
       Parameters          Values
Number of sequencers 5, 10, 15, 20
Batch size for process transactions 500, 1000, 1500, 2000
The ratio of sequencers running SGX, TDX, and CSV 1:2:2
VII-A System Implementation and Setup
We build a prototype of TeeRollup with Golang [19] and develop the smart contracts using Solidity 0.8.0 [20]. We deploy the smart contracts on the Ethereum test network, Sepolia [21], which implements the EVM consistent with the Ethereum main network. We use Intel SGX [14], Intel TDX [15], and Hygon CSV [18] as the TEE platform [23] in a ratio of 1:2:2 for our experiments. We deployed our evaluation on up to 20 nodes, SGX, TDX, and CSV in a ratio of 1:2:2. Specifically, nodes with SGX, TDX, and CSV run on Aliyun ECS g7t.2xlarge with 8vCPU (Intel®Xeon), 8i.2xlarge with 8vCPU (Intel®Xeon), and g7h.2xlarge with 8vCPU (Hygon C86-3G 7390), respectively. We conduct experiments in a LAN environment with 0.5 ± 0.03 ms and a WAN environment with 25 ± 0.1 ms, each machine is equipped with 3GB bandwidth.
We vary parameters in Table III with default values in bold to evaluate system gas cost, throughput, and latency. The experiment focuses on the cost of the chain and the throughput and delay of off-chain transaction processing. Precisely, cost measures the gas cost for executing functions of the TSC. Throughput measures the number of transactions that sequencers can handle per second. Meanwhile, latency measures the time cost for sequencers to vote and process transactions.
VII-B On-Chain Cost of Functions
The on-chain execution costs of TeeRollup are measured in gas, which is a unit that measures the computational effort required to execute operations on Ethereum [22]. The gas consumption depends on the transaction size and execution costs of smart contracts. To make these costs more understandable, we convert gas into USD with a gas price of 20 Gwei ($2\times 10^{-8}$ ETH) and an ETH price of 3655 USD on 10 June 2024.
TABLE IV: Protocol Execution Cost. Execution costs are in USD as per exchange rates of 10 Jun. 2024: ETH/USD 3655.14 and gas price is 20 Gwei ($2\times 10^{-8}$ ETH).
     Methods Cost
Gas USD
Deposit 48551 3.55
UpdateState 156,263 11.42
StartChallenge 47,118 3.44
ResolveChallenge 146,618 10.72
SettleRollup 29,078 2.13
SettleWithdraw 124,511 9.10
Simple ETH transfer 21,000 1.54
Table IV shows our excellent performance in terms of on-chain cost. We measure the complete process of deposit, state update, and challenge on Ethereum when the number of sequencers is 10 (see Sec. V). The UpdateState is one of the most gas-consuming methods, which costs about 156K gas (11.42 USD). That is because it includes the verification of the multi-signature of TEEs and the storage of new state on-chain. However, since 2000 transactions are executed for generating a new state, this fee can be only 78 gas (0.006 USD) per transaction, far less than the gas consumption of a Simple ETH transfer (1.54 USD).
In terms of the challenge, a round of challenge-resolve process consists of the StartChallenge and ResolveChallenge methods, which totally require 194K gas (14.16 USD). The SettleRollup and SettleWithdraw methods, which are crucial for the redemption of the client, totally consuming 154K gas (11.23 USD). However, this only occurs when the TeeRollup needs to be settled. In normal case, the client’s withdrawal only requires a transfer in TeeRollup (see Sec. V-D), which costs 78 gas (0.006 USD).
VII-C Performance Evaluation
We evaluate two performance metrics: 1) throughput, measured in thousands of transactions per second (KTPS), representing the number of transactions executed per second, and 2) latency measured in milliseconds (ms), denoting the average end-to-end delay from the moment sequencers get the transactions until the submission of the transaction.
Throughput and latency. Fig. 4 illustrates the throughput and latency of TeeRollup across varying numbers of sequencers and batch sizes in both LAN and WAN environments. In the LAN environment, throughput shows a modest decrease and latency increases as the number of sequencers rises. This is due to the increase in messaging overhead, causing bandwidth inefficiencies. Throughput exhibits an upward trend with increasing batch size, as larger batch sizes allow TeeRollup to process more transactions concurrently. However, latency also increases with batch size because sequencers need more time to process the transactions in larger batches, resulting in higher delays.
In the WAN environment, the throughput and latency exhibit trends similar to those observed in the LAN environment. However, due to increased communication delay, the WAN performance shows lower throughput and higher latency compared to LAN. Additionally, as batch size increases, the rise in latency in the WAN environment is less pronounced than in LAN. This difference occurs because, in WAN, the primary factor affecting latency shifts from transaction processing speed to communication delay.
(a) Throughput in LAN
(b) Latency in LAN
(c) Throughput in WAN
(d) Latency in WAN
Figure 4: Transaction processing performance with varying different numbers of sequencers and batch size in LAN and WAN.
TEE overhead. Fig. 5 show the influence of introducing TEE on throughput and latency in LAN and WAN, respectively. The operations executed inside a TEE need encryption and to switch context from the regular execution environment of the host machine, which degrades performance. Thus, TEEs’ protection can reduce the throughput and magnify latency. Importantly, even with TEEs’ hardware protection in WAN, the reduction in performance does not exceed the acceptable range for users, maintaining a fine balance between security and performance.
(a) Throughput in LAN
(b) Latency in LAN
(c) Throughput in WAN
(d) Latency in WAN
Figure 5: System performance with/without SEV in LAN and WAN.
VII-D Comparison with Counterparts
We compare TeeRollup with existing rollup schemes, including StarkNet [10], Scroll [28], Optimism [4], and Arbitrum [7]. The average transaction fees for StarkNet and Scroll over the past month were approximately 0.043 USD and 0.114 USD, respectively [6]. The average transaction fees for Optimism and Arbitrum were significantly lower, at 0.012 USD and 0.005 USD, respectively. As demonstrated in Sec. VII-B, the average transaction fee for TeeRollup is 0.006 USD with a package size of 2000 (same as StarkNet), which is reduced 86% than that of the StarkNet. Thus, the transaction fee in TeeRollup is comparable to those of optimistic rollups. As for throughput, current rollup solutions are constrained by the Ethereum mainnet, which can support up to 3,000 TPS. As shown in Sec. VII-C, TeeRollup can achieve throughput exceeding 5000 TPS. Regarding withdrawal times, StarkNet and Scroll offer withdrawals within a few minutes, while Optimism and Arbitrum have a withdrawal period of one week. Although TeeRollup must undergo a challenge process when TEEs are unavailable, the withdrawal time during the normal case remains a few minutes.
VIII Conclusion
In this paper, we propose TeeRollup, a high-performance, cost-effective rollup solution that leverages TEE to streamline the rollup process efficiently. TeeRollup employs a group of sequencers protected by heterogeneous TEEs to process transactions off-chain and DAPs to provide data availability. The cost-efficient challenge mechanism safeguards the redeemability of TeeRollup, even when all TEEs are unavailable. Furthermore, TeeRollup employs the lazy penalty game to incentivize DAPs to work diligently. The prototype implementation of TeeRollup delivers promising results, surpassing existing state-of-the-art approaches in terms of on-chain cost and off-chain processing time.