### v2

# TeeRollup: Efficient Rollup Design Using Heterogeneous TEE

Abstract
Rollups have emerged as a promising approach to improving blockchains’ scalability by offloading transactions execution off-chain. Existing rollup solutions either leverage complex zero-knowledge proofs or optimistically assume execution correctness unless challenged. However, these solutions suffer from high gas costs and significant withdrawal delays, hindering their adoption in decentralized applications. This paper introduces TeeRollup, an efficient rollup protocol that leverages Trusted Execution Environments (TEEs) to achieve both low gas costs and short withdrawal delays. Sequencers (i.e., system participants) execute transactions within TEEs and upload signed execution results to the blockchain with confidential keys of TEEs. Unlike most TEE-assisted blockchain designs, TeeRollup adopts a practical threat model where the integrity and availability of TEEs may be compromised. To address these issues, we first introduce a distributed system of sequencers with heterogeneous TEEs, ensuring system security even if a certain proportion of TEEs are compromised. Second, we propose a challenge mechanism to solve the redeemability issue caused by TEE unavailability. Furthermore, TeeRollup incorporates Data Availability Providers (DAPs) to reduce on-chain storage overhead and uses a laziness penalty mechanism to regulate DAP behavior. We implement a prototype of TeeRollup in Golang, using the Ethereum test network, Sepolia. Our experimental results indicate that TeeRollup outperforms zero-knowledge rollups (ZK-rollups), reducing on-chain verification costs by approximately 86% and withdrawal delays to a few minutes.
Index Terms:
Blockchain, Rollup, Scalability, Trusted Execution Environment
I Introduction
Since the first adoption in Bitcoin [1] in 2008, blockchain technology has experienced remarkable growth across extensive decentralized applications. However, as the number of application requests increases, blockchains have encountered severe congestion due to their low throughput, i.e., scalability issues [2]. To improve the scalability, rollups [3, 4] that offload a substantial portion of transaction execution off-chain emerged as a promising approach. Specifically, they can aggregate the execution result of a batch of off-chain transactions into a single transaction (i.e., a state transition of the rollup) on the blockchain (commonly referred to as the main chain), whereas the main chain only needs to verify the validity of the uploaded state transitions. Due to the promising potentials, rollups have garnered significant attention and research efforts [5]. As of January 2025, statistics from L2BEAT [6] reveal 51 rollup projects in the market, collectively holding a locked value of approximately 57 billion USD. Among these, Arbitrum [7] alone boasts 16.5 million active accounts [8].
Current rollup solutions can be classified into two categories based on verification methods: zero-knowledge rollups (i.e., ZK-rollups) [9, 10, 11] and optimistic rollups (i.e., OP-rollups) [4, 7]. However, these solutions either have high on-chain costs or prolonged withdrawal delays, posing significant challenges to their adoption, particularly in cost-sensitive and time-critical applications. Specifically, ZK-rollups rely on complex zero-knowledge proofs for execution validation on the main chain, but this approach incurs significant gas costs. For instance, verifying a proof generated by ZK-SNARK [12] on Ethereum consumes over 600K gas (approximately \$39 USD). Additionally, generating these proofs is time-intensive and poses challenges for developing, deploying, and supporting smart contracts. In contrast, OP-rollups execute transactions optimistically without immediate verification but face delays in fund withdrawals. This is because OP-rollups introduce a dispute period to ensure security, allowing anyone to detect and report invalid state transitions. For example, systems like Arbitrum and Optimism impose a withdrawal delay of up to one week to maintain security [7, 4].
To break the trade-off between fast withdrawal and low on-chain cost, we propose TeeRollup, which leverages nodes equipped with TEEs to execute transactions off-chain. Specifically, TeeRollup only requires the main chain to verify the TEE-generated signatures. Besides, users can utilize remote attestation during initialization to verify the code running within the TEEs, at which point a key pair is generated inside the TEE. Then, nodes execute transactions within the TEE and generate state transition signed by the TEE, which is verified by the blockchain. Ideally (i.e., without attacks), the integrity provided by TEEs ensures correct execution, preventing malicious nodes from modifying user account status and generating fault state transition. Thus, the main chain can trust the state transition of rollup through simple signature verification.
Unfortunately, directly using TEEs introduces security risks due to their vulnerabilities and unavailability. First, prior research has identified vulnerabilities to compromise TEEs’ integrity and confidentiality [13, 14]. Once a TEE is compromised, a malicious node can disrupt correct transaction execution and generate faulty states with valid signatures from the TEE. To address this issue, TeeRollup adopts a distributed system of nodes with heterogeneous TEEs (e.g., Intel SGX [15], Intel TDX [16], AMD SEV [17], ARM TrustZone [18] and Hygon CSV [19], etc.). In TeeRollup, a state transition must be signed by multiple TEEs to ensure that a single compromised TEE cannot arbitrarily generate a faulty state transition with valid signatures. The states of TeeRollup are organized in a chained structure, where each new state transition is based on the latest state recorded on the main chain. Nodes compete for the right to submit state transitions. However, the main chain will only accept the first valid state transition with sufficient signatures from TEEs, while subsequent submissions are disregarded. Additionally, we utilize on-chain registration and attestation, eliminating the need for mutual attestation between heterogeneous TEEs.
Second, TEEs’ I/O is manipulated by their hosts (i.e., hypervisors or software systems), allowing malicious nodes to drop messages to and from their TEEs. Since the number of nodes in TeeRollup is small (e.g., tens of entities), there may not be enough honest nodes to ensure system availability. If TEE availability is disrupted, users cannot redeem their deposited funds promptly, leading to the locking of user funds on the blockchain. To mitigate this redeemability issue, TeeRollup introduces a challenge mechanism on the blockchain, enabling users to redeem their deposits without relying on sequencers. Specifically, users can initiate challenges on the main chain, and redeem their deposit on the main chain with proof of their balance if no response is received. Here, to generate the proof of balance, the metadata of the system state is required, not just the digest.
Third, to reduce the on-chain storage cost, TeeRollup also incorporates data availability providers (DAPs) to store metadata off-chain. Specifically, TeeRollup only records the digest of metadata on the main chain, and offloads the complete states data and transactions data to the off-chain DAPs. However, the DAPs can be lazy and discard the metadata that they are supposed to store to minimize their costs. Thus, we introduce a collateral scheme to punish dishonest behaviors and a laziness penalty mechanism to incentivize DAPs’ participation.
We build a prototype of TeeRollup with Golang [20], developing the on-chain smart contracts using Solidity 0.8.0 [21] and deploying them on the Ethereum test network, Sepolia [22]. Our experiments use mainstream off-the-shelf TEE devices, including AMD SEV, Intel SGX, and Intel TDX. However, our protocol does not rely on certain platforms and can be extended to more TEEs like ARM CCA once they are mature. These experiments can be easily extended to other virtual machine (VM)-based TEEs.
Contributions. We make the following contributions:
  • $\bullet$
    We introduce TeeRollup, an efficient rollup design that leverages TEEs for off-chain transaction execution. TeeRollup utilizes a group of heterogeneous TEEs, designed to tolerate the compromise of TEEs. Additionally, TeeRollup incorporates data availability providers to reduce on-chain storage costs.
  • $\bullet$
    We address the redeemability issue stemming from TEEs’ unavailability. Our challenge mechanism ensures the redeemability of user funds, even during periods of system crash because of TEEs unavailability. Besides, we implement penalization mechanisms to ensure the diligent operation of data availability providers.
  • $\bullet$
    We conduct experimentation on the prototyped TeeRollup to evaluate its performance compared to ZK-rollup and OP-rollup. Our experimental results demonstrate that the transaction fee in TeeRollup is significantly lower at 0.006 USD compared to StarkNet’s 0.043 USD, showcasing the efficiency of our approach. Moreover, TeeRollup maintains a normal withdrawal time of a few minutes, consistent with StarkNet and better than OP-rollup schemes such as Optimism.
Roadmap. We introduce related work in Sec. II and the system model and goal in Sec. III. The preliminaries are provided in Sec. IV We present the system design in Sec. V, followed by the security analysis in Sec. VI. The system performance is evaluated in Sec. VII. Finally, the paper is concluded in Sec. VIII.
II Background
II-A Revisiting Rollup Solutions
Rollup solutions alleviate the blockchains’ limitations including low throughput, high fees, and network congestion by offloading transactions processing off-chain [23]. Rollup networks are composed of sequencers who aggregate a batch of off-chain transactions into a single state transition. The main chain only needs to verify the validity of the state transition. According to the verification methods, current rollup schemes can be divided into two categories based on their : ZK-rollups and OP-rollups, as introduced below.
ZK-rollup models, as embodied by implementations like StarkEx [9], generate zero-knowledge proofs [24] for the validity of the off-chain state transition. The emergence of solutions like StarkNet [10] and Scroll [11], involve Ethereum Virtual Machine (EVM) compatible zero-knowledge proof. This unique capability extends to the verification of contracts (not just transactions) execution correctness, encompassing both input and output validity in the process. Thomas et al. further introduce an on-demand ZK-rollup creation service [25], which allows several ZK-rollups to co-exist as groups on the same smart contracts, and application examples for Internet of Everything (IoE) [26]. However, the on-chain verification of zero-knowledge proofs in ZK-rollup systems necessitates a notable gas expenditure. For instance, the gas consumption to verify the zero-knowledge proof generated by zk-SNARK on Ethereum is more than 600K gas (about \$39). For zk-STARK, the gas consumption is even higher, reaching approximately 2.5M gas (about \$162) [12].
OP-rollup solutions, implemented by Optimism [4] and Arbitrum [7], pivot towards an optimistic approach for transaction execution. These solutions optimistically assume that transactions are correctly executed by sequencers unless a challenger disputes the execution results. The dispute can be solved on the main chain. The advantage here lies in omitting on-chain verification, reducing the gas cost. Nevertheless, OP-rollups significantly extend the withdrawal time, as they require sufficient time for challengers to verify the proof. For instance, Arbitrum and Optimism enforce a withdrawal period of one week for security [7, 4].
Summary. Existing rollup solutions suffer from high gas costs, low compatibility, and long withdrawal delays. A comparison between existing rollup solutions and TeeRollup (proposed in this paper) is provided in Table I.
TABLE I: Rollup Solution. The Dec, EVM Comp, and Data Avail are short for Decentralized, EVM Compatibility, and Data Availability, respectively. Decentralized property means that the transaction cannot confirmed by a single sequencer. The EVM compatibility denotes the ability to support EVM. For Data Availability, Main Chain and Delegated means the data availability are provided by the main chain and a third party, respectively. Withdrawal time signifies the waiting time for users to retrieve the funds they locked in the main chain.
Rollup Solution Dec EVM Comp Data Avail Withdrawal time
StarkEx [9] Main Chain Few minutes
StarkNet [10] Main Chain Few minutes
Scroll [11] Main Chain Few minutes
Optimism [4] Main Chain One week
Arbitrum [7] Main Chain One week
TeeRollup Delegated Few minutes
II-B Rollup Data Availability
One critical aspect of rollup solutions is ensuring data availability, which refers to the accessibility and integrity of transaction data necessary for maintaining the security and trustworthiness of the system [27]. Storing complete data on the main chain inherits the security and data integrity of the main chain, but faces challenges such as network congestion and high gas costs. For instance, storing a 256-bit integer on the Ethereum smart contract STORAGE field costs 8K gas (0.52 USD) [21]. Therefore, it is unpractical to simply permanently store the complete data of the rollup on the main chain. StarkEx [9] and StarkNet [10] address data availability issues by including aggregated compressed transactions in the CALLDATA field of the main chain, which reduces the gas cost [28]. However, this method does not allow for permanent storage of transactions and consumes more gas compared to only recording the digest of state on the main chain. Thus, we choose to only record the digest of rollup state (i.e., the state) on the main chain with complete data (i.e., metadata) stored off-chain by the data availability providers (DAPs).
II-C Trusted Execution Environment
Trusted Execution Environments (TEEs) aim to run applications in a secure environment without leaking secrets to an adversary who controls the computing infrastructures. Specifically, TEEs provide enclave to run code and the attestation mechanism to prove the correctness of computation. Influential TEE implementations include Intel SGX [15], Intel TDX [16], ARM TrustZone [18], ARM CCA [29], AMD SEV [30] and Hygon CSV [19].
Recently, TEEs have been widely used in blockchain designs to enhance security, privacy, and performance [31, 32, 33, 34, 35, 36]. Teechain [31] establishes a payment platform under TEE protection, while Tesseract [37] and Bool Network [38] employs TEE to ensure the security of cross chain transactions. Tommaso et al. [32] and Xu et al. [33] leverage the TEE for the correctness and privacy of off-chain execution, respectively. These works assume that TEEs provide integrity and confidentiality guarantees, where the adversary cannot know the private key inside the TEE.
However, many studies indicate that TEEs are vulnerable to various attacks such as side-channel attacks [39], unprotected I/O [40], and ASID abuses [41]. Thus, in our work, we assume TEEs can be compromised, i.e., no integrity and confidentiality properties, which differs from prior work with the perfect assumption of TEEs [31, 32, 33, 34, 35, 36]. In particular, the adversary can steal the private keys for signing messages from the compromised TEE. To tolerate malicious behaviors of compromised TEEs, one approach is using a distributed system of heterogeneous TEEs. This heterogeneity can arise from using TEEs from different vendors, different types of processors, or different levels of security requirements [42]. Since different TEE platforms adopt different hardware and software architectures, it is difficult to breach their security simultaneously. Besides, when a TEE is compromised, manufacturers will timely resolve these issues, making security compromise of all TEEs nearly impossible [43, 44].
III Problem Statement and System Model
In this section, we first present the system model and threat model. Then, we introduce the associated challenges and solutions. Table II lists the frequently used notations in this paper.
TABLE II: Summary of notations.
Term Description Term Description
$\mathcal{M}$ Main chain $st_{h}$ State of TeeRollup
$h$ Height of $st_{h}$ TSC TeeRollup smart contract
MSC Manager smart contract $n$ Number of sequencers
TToken Tokens of TeeRollup $p_{i}$ Sequencers in TeeRollup
$T_{addr}$ TSC account on $\mathcal{M}$ $\eta_{i}$ Enclave for sequencer $p_{i}$
$(pk_{i},sk_{i})$ Key pair for enclave $i$ $\tau$ Timer for resolve challenge
III-A System Model
We consider a system of $n$ sequencers, denoted by the set $\mathcal{G}=\{p_{1},p_{2},...,p_{n}\}$. We follow the existing assumption [45], where each sequencer $p_{i}$ is equipped with one type of TEE platform (e.g., Intel SGX, Intel TDX, etc.), denoted by $\eta_{i}$. We assume at most $f$ sequencers’ TEEs are compromised at any time, leaving the remaining $n-f$ sequencers’ TEEs uncompromised. Here, since TEE machines adopt different architectures and are produced by different manufacturers, it is difficult to breach the security of $f$ ones at the same time. The parameter $f$ can be determined by the rollup service provider. There is a public/private key pair of the sequencer $p_{i}$, denoted by $(pk_{i},sk_{i})$, in which the private key is only stored and used inside TEE.
III-B Threat Model
We model the malicious sequencers as Byzantine adversaries, i.e., they can behave in arbitrary ways. The threat model of sequencers consists of three cases: the honest sequencers with/without compromised TEEs, the malicious sequencers with compromised TEEs, and the malicious sequencers with uncompromised TEEs.
  • $\bullet$
    Case 1: Honest sequencers with uncompromised TEEs. The honest sequencers with their TEEs do not deviate from the protocol.
  • $\bullet$
    Case 2: Honest sequencers with compromised TEEs. The honest sequencers with their TEEs do not deviate from the protocol.
  • $\bullet$
    Case 3: Malicious sequencers with uncompromised TEEs. The malicious sequencers have full control over the operating system of their TEEs, including root access and control over the network. The sequencers can arbitrarily launch, suspend, resume, terminate, and crash TEEs at any time. Besides, the sequencers can delay, replay, drop, and inspect the messages sent to and from TEEs, i.e., manipulating input/output messages of TEEs. In other words, the TEE cannot guarantee availability due to these manipulations.
  • $\bullet$
    Case 4: Malicious sequencers with compromised TEEs. The sequencers have full control of the enclave and know the private key $sk_{i}$ inside the compromised TEE. In this case, TEEs have no integrity and confidentiality guarantees. We assume static corruption by the adversary, where a fixed fraction of all sequencers’ TEEs is compromised.
Existing TEE-based blockchain research [32, 37] only focuses on case 1 and case 3, without addressing the scenario of compromised TEEs. We extend this by considering up to $f$ compromised TEEs of sequencers and scenarios where all sequencers can be malicious. For data availability providers, we follow the existing assumption in [46] that they are all rational.
Main chain model. We assume the TeeRollup is built on the main chain $\mathcal{M}$. The main chain provides finality (i.e., once a transaction is included in a block it is considered final) [47] and enables smart contracts [48]. The finality of a transaction can be verified by checking whether the including block satisfies certain public rules [49] (i.e., the block is correctly formatted, contains valid signatures, and so on). Smart contracts can access the current time using the method $block.timestamp$ and provide cryptographic schemes including hash computing and ECDSA encryption [21].
III-C System Goals
The rollup system issues tokens (denoted by TTokens), and allows clients to transfer tokens in the rollup or redeem tokens on the main chain. A client on the main chain $\mathcal{M}$ with account $c_{addr}$ can enter rollup by depositing on the account $T_{addr}$ held by the rollup. Then, the rollup issues the TTokens and locks the deposit on the account $T_{addr}$. Clients can submit transactions to the sequencers to transfer within the account of the rollup. Sequencers handle transactions and update the state of the rollup to the blockchain $\mathcal{M}$.
The state of the rollup can be denoted as $st_{h}$, where $h$ is the sequence number (referred to as the height) of a state $st$. The processing of transactions inside rollup can be defined as $st_{h+1}\leftarrow execute(st_{h},txs_{h+1})$, where $txs_{h+1}$ is a batch of transactions. The rollup system has to satisfy the following two properties:
  • $\bullet$
    Correctness: TeeRollup must ensure the integrity and correctness of state transitions on the main chain.
  • $\bullet$
    Redeemability: Any client can redeem their TTokens in TeeRollup.
III-D Design Challenges
To better outline the TeeRollup design, we first introduce a strawman solution (referred to as SRollup) that leverages TEE to shield the sequencer from malicious behavior. In SRollup, a single sequencer collects and executes the transactions from clients and submits the execution result to the main chain. While SRollup always provides the fundamental functions of issuing, transferring, and redeeming TTokens, it does not achieve the Redeemability defined in III-C. This is because that SRollup does not consider the compromised TEE and potential malice of sequencers and data availability providers.
Compromised TEE. As mentioned in  Sec. III-B, TEEs can be compromised, allowing malicious sequencers to acquire the secret keys of their compromised TEEs. Malicious sequencers with compromised TEEs can forge false transactions or execute transactions incorrectly to generate incorrect states with valid signatures. Consequently, a single TEE cannot be fully trusted by the system. To mitigate TEE compromises, TeeRollup employs a group of sequencers equipped with independent and heterogeneous TEEs. Sequencers validate and sign state transitions within their TEEs, competing for the right to submit them. The main chain will only accept the first valid state transition, based on the latest state, along with sufficient signatures. Additionally, TeeRollup uses on-chain attestation during registration, avoiding direct attestation between heterogeneous TEEs.
Malicious sequencers. According to our threat model (see Sec. III-B), the operating system of the TEE is fully controlled by the sequencer. Malicious sequencers can crash the TEE, censor transactions, or selectively filter out certain client requests. If there are insufficient available TEEs, TeeRollup cannot process client requests, preventing clients from redeeming their deposits and compromising redeemability. To address this, TeeRollup introduces a challenge mechanism to ensure redeemability and enable on-chain settlement when TEEs fail to provide service. If a client finds their transaction unprocessed for an extended period, they can submit the transaction data to the main chain via a challenge method implemented in the smart contract. When a challenge is initiated, a timer starts. If the sequencer fails to respond within the allotted time, the system is deemed unavailable and all the deposit will be settled on the main chain. This mechanism incentivizes sequencers to respond promptly to challenges, preventing system-wide settlement while ensuring clients can redeem their funds even during system unavailability.
Lazy data availability providers. The states of TeeRollup are recorded on the main chain, which guarantees its integrity (see Sec. III). Therefore, the correctness of the metadata provided by DAPs can be verified by the $state$ recorded on smart contract (i.e., DAPs cannot forge account balance, transaction data, or other metadata). However, the DAPs can show laziness, discarding the data they are supposed to store. Thus, clients and sequencers cannot get metadata of the $state$ on smart contracts. TeeRollup introduces collateral and laziness penalty mechanisms to incentivize DAPs to behave diligently. DAPs need to provide the collateral when entering the network and once a DAP acts dishonestly, some or all of its collateral will be deducted.
IV Preliminaries
In the following, we introduce the cryptography primitives, Merkle tree, and the model of TEEs used in the TeeRollup protocol.
[…]
Merkle tree. Merkle tree is a binary tree data structure in which each leaf node contains the hash of a data block (e.g., the key-value pair), and each non-leaf node contains the hash of its child nodes. Constructed bottom-up, a single hash at the root can be generated, representing the entire dataset’s integrity. The root of a Merkle Tree is integrated into each block, serving as a comprehensive state digest, enabling sequencers to maintain and verify the state digest efficiently. The Merkle Tree allows leaves to generate a Merkle path for verifying specific key-value pairs under a root, denoted as $\delta\leftarrow root.proof(\left<k,v\right>)$. Anyone can verify this pair of key-value through $verifyproof(\delta,\left<k,v\right>)$.
Figure 1: An overview of TeeRollup architecture. The red components represent smart contracts on the main chain, the blue components indicate operations executed within the TEE, and the gray components correspond to the storage provided by data availability providers.
Trusted Execution Environments (TEEs). The sequencers can instruct their TEEs to create new enclaves, i.e., new execution environments running a specific program. We follow prior work [50] to model the functionality of TEEs.
  • $\bullet$
    $eid\leftarrow install(prog)$: Installing the program $prog$ as a new enclave with a unique enclave ID, $eid$ within the TEE.
  • $\bullet$
    $(outp,\sigma)\leftarrow resume(eid,inp)$: Resuming the enclave from the crash with ID $eid$ to execute program $prog$ using the input $inp$. $\sigma$ is the TEE’s endorsement, confirming that $outp$ is the valid output.
For uncompromised TEE, the verification of the output $outp$ is trusted by users when $verify(\sigma,eid,prog,outp)=1$, indicating the successful attestation of the TEE. Besides, we assume an attestation API for TEE to generate an attestation quote $\rho\leftarrow attest(eid,prog)$ proving that the program $prog$ has been installed in the enclave $eid$. And $\rho$ can be verified through method $verifyquote(\rho)$.
V TeeRollup Design
V-A Overview
Fig. 1 shows the architecture of TeeRollup, a rollup based on TEE. TeeRollup considers four roles: the TeeRollup Smart Contract (TSC) responsible for record-keeping on the main chain, the Manager Smart Contract (MSC) for managing sequencers, the TEE-equipped sequencer committee processing off-chain transactions, and the data availability provider (DAP) storing metadata for TeeRollup, as shown in Fig. 2.
  • $\bullet$
    TeeRollup smart contract (TSC). TSC records the TeeRollup state proof in the form of a state (rather than complete metadata) to reduce on-chain storage cost.
  • $\bullet$
    Manager smart contract (MSC). MSC manages registered TEEs of sequencers, especially assigning an ID for them and recording the private key $m_{pk}$ inside TEE.
  • $\bullet$
    Sequencer. Sequencer primarily collects and executes transactions from clients, submits the state to TSC, and sends metadata to DAP.
  • $\bullet$
    Data availability provider (DAP). DAP accepts and stores the metadata sent by sequencers to ensure access to system metadata at any time. The metadata includes the state of the system (i.e., account balances), the update history, and transaction data.
There are four key procedures in TeeRollup. First, during the initialization phase, the service provider deploys both the TSC and the MSC on the blockchain $\mathcal{M}$. Subsequently, sequencers register their TEEs on the MSC (see Sec. V-C). Second, clients deposit funds on the TSC and request the issuance of TTokens in TeeRollup, which can be used for transfers. Sequencers receive transactions from clients, process these transactions, and generate an updated state. Then, the sequencers broadcast the new state along with metadata and collect votes. Once enough votes are collected, the sequencers submit the state of TeeRollup to the TSC and send the metadata to the DAPs (see Sec. V-D). Third, if clients’ transactions are not processed in time, clients can challenge the TSC to enforce the transaction. Even if the entire system cannot respond, they can get a refund directly from the TSC to avoid the property loss (see Sec. V-E). Fourth, to better regulate the behavior of DAPs and prevent laziness, we use the incentive method of collateral plus laziness punishment (see Sec. V-F).
V-B Data Structure
State format. In TeeRollup, the state acts as a digest of the metadata and anyone can verify the transactions, balance, and other information of TeeRollup using this digest. The state $st_{h}$ on height $h$ has the following structure:
$st_{h}:=\left<h,H(st_{h-1}),R_{h},H(txs_{h})\right>$
where $H(st_{h-1})$ refers to the hash of the previous state $st_{h-1}$, $R_{h}$ is the Merkle root of the account tree of TeeRollup, and $H(txs_{h})$ is the hash of the transactions list executed to generate the $st_{h}$. Account tree is a Merkle Tree that stores the clients’ account in TeeRollup, denoted as $A_{h}$. Account tree $A_{h}$ consists of the account address and the balance for all accounts.
In TeeRollup, the states follow a chain structure, which has been adopted by Bitcoin [1], Ethereum [49], and some BFT protocols [51]. Every state contains the hash of the previous state, and the state can be indexed by height (i.e., the distance from the initial state). If there is a path from $st_{h+1}$ to $st_{h}$ then we say that $st_{h+1}$ extends $st_{h}$ In TeeRollup, there is only one state at each height. Once the latest state is obtained, the correctness of any historical state can also be verified.
Vote and certificate. To achieve collaboration among multiple sequencers, TeeRollup allows sequencers to compete for the right to submit states. Specifically, any sequencer can become the leader, broadcast the state, and collect votes. However, the TSC will only accept the first state that arrives at one height. A vote $v_{i}$ from the sequencer $p_{i}$ of state $st_{h}$ has the following structure:
$v_{i}:=\left<H(st_{h}),pk_{i},\sigma_{i}\right>$
where $H(st_{h})$ is the hash of $st_{h}$, $pk_{i}$ is the public key for sequencer $p_{i}$, and $\sigma_{i}$ is a signature created by the sequencer $p_{i}$ over $H(st_{h})$. Here, the signature is generated inside the TEE of sequencer $p_{i}$ with the private key $sk_{i}$. If there is a set of $f+1$ signatures, it forms the quorum certificate (QC) for the state. Here, a QC can be implemented as multi-signatures or aggregated signatures.
Figure 2: An overview of the main functions provided by TeeRollup.
V-C Sequencers Registration and Configuration
TeeRollup implements the MSC to manage the sequencer committee, which provides the registration and configuration for sequencers.
Enclave registration. Sequencers equipped with TEEs can participate in TeeRollup by creating an enclave $\eta_{i}$ and registering it on the MSC. To clarify the process, we present an example of a sequencer $p_{i}$ registering on the MSC for clarity. First, the sequencer $p_{i}$ creates an enclave $\eta_{i}$ and installs the TeeRollup program $prog$ inside it. Upon the creation of $\eta_{i}$, an asymmetric key pair $(pk_{i},sk_{i})$ is generated. Notably, the public key $pk_{i}$ also serves as an account address for $\eta_{i}$ on the main chain [49]. The secret key $sk_{i}$ is securely stored within $\eta_{i}$, and the public key $pk_{i}$ is returned as output to the sequencer $p_{i}$.
Then, the sequencer $p_{i}$ registers $\eta_{i}$ by invoking Register with $\eta_{i},\rho_{i},pk_{i}\right>$ on the MSC. Here, $\rho_{i}$ is a proof generated by $\eta_{i}$, asserting its execution of the program $prog$. The TSC verifies that $verifyquote(\rho_{i})=1$ [50, 52]. Upon successful verification, the MSC adds the sequencer $p_{i}$ and $pk_{i}$ to the sequencer list. The registration process guarantees the correct loading of the program $prog$ for all registered enclaves and ensures that the secret key $sk_{i}$ remains confidential during registration. Since the heterogeneous TEEs can be registered and attested on the MSC, they can rely on the MSC to establish trust without needing mutual attestation. As a result, sequencers with heterogeneous TEEs can trust each other and use key pair $pk_{i},sk_{i}\right>$ to generate signatures between TEEs. There is no need to re-attest enclaves in subsequent protocol steps.
Sequencer configuration. As described earlier, sequencers register their enclaves on the MSC, where their identities and public keys are securely recorded. The MSC manages the sequencers and provides authentication services to the TSC. Specifically, the MSC maintains a list of sequencers, including their public keys and the committee strategy used by the TSC. When the TSC needs to verify messages sent by sequencers, it can invoke the VerifyQC function of the MSC to ensure the authenticity of the QC. In this paper, we adopt a voting strategy, where a QC is considered valid if it includes signatures from more than $f+1$ sequencers in the committee.
V-D Normal-Case Operations
As shown in Fig. 2, TeeRollup provides issue, transfer, and redeem functions for clients.
V-D1 Issue
Client deposits in the TSC and the TeeRollup issues the corresponding amount of TTokens for the client. The exchange rate between the currency of blockchain $\mathcal{M}$ and TToken is beyond the scope of this paper. The rate is set to be 1 for simplicity.
Step 1. The client invokes the Deposit on the TSC, denoted as a tuple of $s_{addr},v\right>_{\sigma}$, where $s_{addr}$ is the client’s address and $v$ is the deposit value (greater than zero), and $\sigma$ is the signature of the client. In TeeRollup, we follow the existing practices in rollup systems, where rollup shares the same key algorithm as the main chain [3]. This means that a client holding an account address (i.e., public key) and its corresponding secret key can use the same key pair on both the main chain and the rollup.
Step 2. The enclave $\eta_{i}$ within the sequencer $p_{i}$ monitors the $deposit$ transaction on TSC. Enclave $\eta_{i}$ issues TTokens and changes the balance of account $s_{addr}$ in TeeRollup. Enclave $\eta_{i}$ batches multiple transactions together to optimize efficiency and executes these transactions, generating the new state $st_{h+1}$ and a $lock$ transaction denoted as $id\right>$, where $id$ is the number of the $deposit$. The sequencer $p_{i}$ sends the $st_{h+1}$ and $lock$ with the metadata to other sequencers, who vote for the $st_{h+1}$.
Step 3. After collecting $f+1$ votes, the sequencer $p_{i}$ generates the QC for $st_{h+1}$ and sends $st_{h+1}$ and the QC to the TSC. In TeeRollup, state $st_{h+1}$ that meets the following conditions can be certificated by the TSC. First, the height of the lasted state in TSC is $h$. Second, $st_{h+1}$ contains the hash value of $st_{h}$, which corresponds to the lasted state in TSC. Third, the QC for $st_{h+1}$ is valid. Once certified, the TSC records $st_{h+1}$ on the main chain. Additionally, the TSC locks the client’s $deposit$, which cannot be withdrawn unless the client exits TeeRollup. The update of $st_{h+1}$ and the locking of the $deposit$ are performed in a single operation on the TSC to ensure atomicity.
step 4. After $st_{h+1}$ and the lock transaction is confirmed on the main chain, the sequencer $p_{i}$ returns the execution result to the clients and sends the metadata to the DAPs.
V-D2 Transfer
Clients transfer their TTokens by sending transactions to the sequencer.
Step 1. A client submits a transaction $tx$ to the sequencers. $tx$ can be denoted by tuple of $s_{addr},r_{addr},v\right>_{\sigma}$, where $s_{addr}$ is the sender’s address, $r_{addr}$ is the receiver’s address, $v$ is the transferred value (greater than zero), and $\sigma$ is the signature of the sender.
Step 2. The execution of $tx$ changes the balance of account $s_{addr}$ and $r_{addr}$. Transaction $tx$ is batched with other transactions into $txs_{h+1}$. Enclave $\eta_{i}$ executes $txs_{h+1}$, and generates the new state $st_{h+1}$ signed with its private key $sk_{i}$. Then, the sequencer $p_{i}$ broadcasts $st_{h+1}$ along with the transactions $txs_{h+1}$ to others and waits for their votes.
Step 3. After collecting $f+1$ votes, the sequencer $p_{i}$ generates the QC for $st_{h+1}$ and sends $st_{h+1}$ and the QC to the TSC. TSC verifies that QC is valid, and then records $st_{h+1}$.
Step 4. If $st_{h+1}$ is confirmed on the main chain, the sequencer $p_{i}$ returns the execution result to the clients and sends the metadata to the DAPs.
V-D3 Redeem
Client burn TTokens in TeeRollup and TSC refund to the Client. To burn TTokens, clients transfer them to the designated account $a_{0}$, which is a burn-only account. Account $a_{0}$ can only receive tokens and cannot transfer them, ensuring that tokens sent to $a_{0}$ are permanently considered burned. TeeRollup employs an off-chain redemption method to minimize gas costs. To address redeemability challenges, TeeRollup incorporates a challenge mechanism, detailed in Sec. V-E.
Step 1. Clients submit requests to the sequencer $p_{i}$, and the request can be denoted as $s_{addr},a_{0},v\right>_{\sigma}$, where $s_{addr}$ is the client’s address, $v$ is the transferred value (greater than zero), and $\sigma$ is the signature of the client.
Step 2. Enclave $\eta_{i}$ batches $tx$ with other transactions, executes them, and generates the new state $st_{h+1}$ with a $refund$ transaction for the TSC, denoted as $s_{addr},v\right>$. Enclave $\eta_{i}$ signs $st_{h+1}$ and $refund$ with its private key $sk_{i}$. Then, the sequencer $p_{i}$ sends $st_{h+1}$ with $refund$ to others and waits for their votes.
Step 3. After collecting $f+1$ votes, the sequencer $p_{i}$ generates the QC for $st_{h+1}$ and $refund$. Then, the sequencer $p_{i}$ submits $st_{h+1}$ and $refund$ along with the QC to the TSC. If the QC is verified as valid, the TSC records $st_{h+1}$ on the main chain and refunds currency with a value of $v$ to $s_{addr}$.
Step 4 After $st_{h+1}$ and $refund$ transaction are confirmed on the main chain, the sequencer $p_{i}$ returns the execution result to the clients and sends the metadata to the DAPs.
1:Function Init
2:    ContractState $\leftarrow$ Active
3:    $Chal\leftarrow\emptyset$
4:    $Dep\leftarrow\emptyset$
5:
6:Function Deposit ($M_{addr},T_{addr},v$)
7:    Require ContractState $=$ Active
8:               $\wedge$ $value>0$
9:    $id\leftarrow$ H($M_{addr},block.timestamp$)
10:    $Dep_{id}.sender\leftarrow M_{addr}$
11:    $Dep_{id}.value\leftarrow v$
12:    $Dep_{id}.solved\leftarrow false$
13:    Trigger Deposit event
14:
15:Function UpdateState($QC,S$)
16:    Require ContractState $=$ Active
17:               $\wedge$ $S.h=height$
18:               $\wedge~{}S.preHash=H(st_{h}eight)$
19:               $\wedge$ VerifyQC($QC$)
20:    $height\leftarrow t+1$
21:    $st_{height}.root\leftarrow S$
22:
23:Function StartChallenge ($M_{addr},tx,pledge$)
24:    Require ContractState $=$ Active
25:               $\wedge$ $pledge\geq Pledge_{C}$
26:    $id\leftarrow$ H($M_{addr},tx,block.timestamp$)
27:    $Chal_{id}.tx\leftarrow tx$
28:    $Chal_{id}.startChal\leftarrow block.timestamp$
29:    Trigger Challenge event
30:
31:Function ResloveChallenge ($id,QC$)
32:    Require ContractState $=$ Active
33:               $\wedge$ VerifyQC($QC$)
34:    Delete($Chal_{id}$)
35:
36:Function SettleRollup ($id$)
37:    Require ContractState $=$ Active
38:               $\wedge$ $block.time-Chal_{id}.startChal>T_{w}$
39:    ContractState $\leftarrow$ Frozen
40:    Trigger Settle event
41:
42:Function SettleWithdraw ($T_{addr},M_{addr},b,\delta,\sigma$)
43:    Require ContractState $=$ Frozen
44:               $\wedge$ Verify($T_{sddr},\sigma$)$=1$
45:               $\wedge$ VerifyProof($\delta,b,T_{sddr}$)$=1$
46:    Refund($M_{addr},b$)
Algorithm 1 Challenge Algorithm of TRollup Smart Contract
V-E Challenge Mechanism
To address the redemption issue, we propose a challenge mechanism deployed in TSC. In TeeRollup, clients can initiate a challenge in TSC if their request is not processed. If there is no response from the sequencers, clients can also withdraw their deposit from TSC with the proof of balance provided by DAP.
First, clients invoke the StartChallenge to initiate a challenge on TSC and receive an $id$ (calculated by the hash of the sender address, transaction, and block timestamp) as a return (cf. Algorithm 1 line 21). Then, TSC sets a timer and triggers an $challenge$ event with $tx,block.timestamp\right>$. If the enclave receives the $challenge$, it executes the $tx$ in the next batch and votes for it before $\tau$ passes (cf. Algorithm 1 line 26-28). Therefore, when the sequencers provide the QC for the $tx$, it proves that the enclaves have received the transaction and executed it. To give the sequencers sufficient time to process the transactions and respond, $\tau$ is set to a duration of several hours. The waiting time $\tau$ for the challenge does not affect the execution efficiency in normal cases.
On the other hand, if the sequencer drops the request, and cannot generate the response to the challenge in time, clients can invoke the SettleRollup($id$) to settle TSC. TSC verifies whether the current block time has exceeded the start time of the challenge plus the waiting time $\tau$. If the verification is passed, the contract state of TSC becomes frozen (cf. Algorithm 1 line 32). At this time, sequencers can no longer update the state, and the deposit is also not allowed, while the client can withdraw their balance. To prevent the client’s malicious challenge (for example, they do not submit a transaction $tx$ to the sequencer, but carry out a challenge on the chain), the client needs to lock collateral on TSC when initiating a challenge (cf. Algorithm 1 line 20). If the challenge fails, the collateral will be forfeited. Additionally, to prevent malicious sequencers from censoring the client’s transaction by dropping messages sent to the enclave, the client must submit the $tx$ to the TSC. This ensures that the transaction is publicly available and can be processed by the sequencers. If the sequencers fail to process the transaction and instead attempt to suppress it by withholding or dropping the transaction, the sequencers will be penalized and rollup will also be settled. If a client finds that its transactions are consistently delayed and require challenges, it can redeem its deposit from TeeRollup.
To withdraw the balance $b$ in account $a$, the client needs to provide two proofs: (i) the client has the key of account a; and (ii) the balance in account a is b. First, according to Sec. V-D, the account address $a$ is the public key and the client can generate the signature with the private key, proving their control of account $a$. Second, the client can get the proof of balance from DAP, i.e., the Merkle proof for the balance of the client’s account. The state recorded in TSC contains the $AccountRoot$, which is the Merkle root of the account balance. Thus, the DAPs can provide the balance $b$ and proof $\delta\leftarrow st_{h}.AccountRoot.proof(a,b)$, where $a$ is the address for the account $b$ in TeeRollup. Client generates the signature $\sigma$ for $\delta$ with the secret key of account $b$ and invoke SettleWithdraw($T_{addr},M_{addr},b,\delta,\sigma$). Finally, TSC verifies the proof of balance and the signature $\delta$. If successful, TSC refund the balance of account $b$ to the client. Thus, every client can redeem their deposit on TSC.
The above design solves the problem that the client’s transfer and withdraw (i.e., a transfer to a specific account) transactions are not executed, but the client’s deposit is also at risk of not being processed. Thus, in our design, every $deposit$ should be confirmed by the sequencers. When clients raise a TTokens request and lock the $deposit$ on the TSC, the $deposit$ is unsolved, and a timer with a countdown of $\tau$ is set. Then, sequencers process the $deposit$, issue TTokens for the client, and submit the new state to TSC. Upon the new state is confirmed on the main chain, sequencers respond to the $deposit$, and set it to be solved. If the $deposit$ is unsolved until $\tau$ passes, TSC can refund $deposit$ to the clients.
V-F Data Availability Provider
To reduce on-chain storage costs, we adopt an off-chain method for ensuring data availability. For a state $st_{h}$, DAPs store the transaction lists $txs_{h}$ and account tree $A_{h}$, which together constitute the metadata of TeeRollup. Alongside the submission of updated states to the TSC, sequencers simultaneously upload the metadata to the DAPs. DAPs store the metadata and provide it to clients when settlement occurs (as mentioned in Sec. V-E). However, DAPs can act lazy (e.g., withholding data or lazy validation) and pretend as if metadata was stored. Therefore, to motivate DAPs to store metadata diligently, we have designed a laziness penalty mechanism to punish the lazy DAP.
Registration. To ensure that DAPs store the metadata of the state recorded on TSC, we require DAPs to register and lock a specified collateral amount. For clarity, we provide an illustrative example of a DAP registering on the TSC. Initially, the DAP registers by submitting $D_{addr},v\right>_{\sigma}$ to TSC, where $D_{addr}$ represents the DAP’s address (i.e., the public key), and $v$ is the collateral value. Upon receipt, TSC authenticates the signature $\sigma$ and validates whether $v$ exceeds the minimum requirement value for collateral, denoted as $C$. If a DAP successfully passes the verification, it is authorized to store the metadata of TeeRollup.
Laziness penalty mechanism. We introduce the collateral $C$ and laziness penalty mechanism to incentivize DAPs. Specifically, the client can initiate random data requests to the DAPs, and the node that fails to post a response loses part or all of its collateral (i.e., a slashing of collateral). Once the data request is triggered, the DAPs should respond to the TSC within time $l$, otherwise, the DAPs will suffer a slashing of collateral. The response consists of the metadata and the Merkle path proof for the metadata.
For the slashing mechanism, we follow the design in [46], which introduced an optimal slashing function for DAPs. In our design, for a system with $m$ DAPs, each DAP stores a copy of the metadata. Thus, any DAP can respond to the request and provide the complete metadata. Besides, the response from DAP can be verified by the TSC, and the invalid response is considered to be a no-response. For a specific data request, the sets $X=\left<x_{1},x_{2},...,x_{m}\right>$ represent the reply of DAP. Let $x_{i}=1$, if DAP $q_{i}$ sends a valid data response to TSC, and $x_{i}=0$ otherwise. Since this design is a special case for [46], the slashing function for $q_{i}$ is defined as follows.
$f_{i}(x)=\begin{cases}0,&i=1\\ -C,&\sum_{j=1}^{m}x_{j}<1~{}and~{}x_{i}=0\\ -W-\epsilon,&\sum_{j=1}^{m}x_{j}\geq 1~{}and~{}x_{i}=0\\ \end{cases}$
The TSC slashes the collateral $C$ put up by each DAP that has not sent a valid response before the timeout if there is no valid response. Otherwise, if there is more than one response in the TSC before the timeout, it punishes the non-responsive DAPs by a modest amount, namely $W+\epsilon$, where $W$ is the cost for the DAP to construct and send a response to TSC. Note that $\epsilon$ is a small number to prevent the DAP from responding negatively to save the cost $W$.
VI Security Analysis
We analyze the security of TeeRollup, which satisfies the correctness, and Redeemability properties.
Lemma 1.
If a malicious sequencer forges an invalid current state $st_{h}^{\prime}\neq st_{h}$, the new state $st_{h+1}^{\prime}$ generated by the enclave will not be accepted by TSC.
Proof.
According to the Sec. V-D, the honest enclave generates the $st_{h+1}^{\prime}$ containing the hash of $st_{h}^{\prime}$, i.e., $st_{h+1}^{\prime}.preHash\leftarrow H(st_{h}^{\prime})$. Thus, if the sequencer submits the $st_{h+1}^{\prime}$ signed by the enclave, TSC will check whether $st_{h+1}^{\prime}.preHash$ is equal to $H(st_{h})$. Obviously, since $st_{h}^{\prime}\neq st_{h}$, TSC will not accept the $st_{h+1}^{\prime}$. ∎
Lemma 2.
If a malicious sequencer with the compromised TEE forges an invalid state $st_{h+1}^{\prime}$ with transaction list $txs_{h+1}^{\prime}$ (i.e., $st_{h+1}^{\prime}$ cannot be calculated by executing transactions $txs_{h+1}^{\prime}$ with the initial state $st_{h}$), TSC will not accept the state.
Proof.
By Lemma 1, the sequencer with the protected TEE cannot make the enclave output a forged state by providing the forged input. If the sequencer provides the correct input $st_{h}$, the enclave will only generate $st_{h+1}$ by executing $txs_{h+1}^{\prime}$. This is because the program in the enclave is protected from execution. Thus, the malicious sequencer with protected TEE cannot generate the forged state $st_{h+1}^{\prime}$.
According to Sec. III, TeeRollup have $n$ sequencers, and up to $f$ TEE can be compromised. Thus, the $st_{h+1}^{\prime}$ can only be signed by at most $f$ compromised enclaves of sequencers, i.e., it will not be accepted by TSC. ∎
Theorem 1 (Correctness).
Malicious sequencers cannot upload an invalid state accepted by the TSC.
Proof.
According to Lemma 1 and Lemma 2, malicious sequencers with uncompromised TEE or compromised TEE cannot forge invalid states accepted by the TSC. ∎
Theorem 2 (Redeemability).
If a client submits a transaction $tx$ to sequencers, the transaction will finally be executed or the TSC will be settled (i.e., the deposit of all clients can be redeemed in the main chain).
Proof.
According to the challenge mechanism, the client can initiate a challenge, if its transaction is not executed for a long time. Once a challenge is initiated, there are two cases as follows.
  • $\bullet$
    Case 1. Sequencers resolve the challenge before the waiting time $\tau$ passes. In this case, at least $f+1$ enclaves of sequencers have received the transaction $tx$. Thus, at least one protected enclave will batch the transaction and execute it to generate a new state.
  • $\bullet$
    Case 2. Sequencers fail to resolve the challenge before the waiting time $\tau$ passes. As a result, the client can invoke the Settle($id$) to settle the TSC. TSC is frozen immediately, and any client can redeem their deposit on TSC.
VII Performance Evaluation
We evaluate TeeRollup in terms of its on-chain verification costs, throughput, and latency of processing off-chain transactions. We consider two state-of-the-art counterparts, StarkNet [10], Scroll [11], Optimism [4], and Arbitrum [7]. Through various experiments, we answer the following questions:
  • $\bullet$
    Q1: How does TeeRollup perform in the on-chain cost?
  • $\bullet$
    Q2: How does TeeRollup perform in throughput and latency?
  • $\bullet$
    Q3: How does TeeRollup perform compared to StarkNet, Scroll, Optimism, and Arbitrum?
VII-A System Implementation and Setup
We build a prototype of TeeRollup with Golang [20] and develop the smart contracts using Solidity 0.8.0 [21]. We deploy the smart contracts on the Ethereum test network, Sepolia [22]. We use Intel SGX [15], Intel TDX [16], and Hygon CSV [19]. We deployed our evaluation on up to 20 nodes, with the distribution of SGX, TDX, and CSV distributed in a ratio of $1:2:2$. Specifically, nodes with SGX, TDX, and CSV run on Aliyun ECS g7t.2xlarge with 8vCPU (Intel®Xeon), 8i.2xlarge with 8vCPU (Intel®Xeon), and g7h.2xlarge with 8vCPU (Hygon C86-3G 7390), respectively. We consider two deployment scenarios: local area network (LAN) and wide area network (WAN). For WAN experiments, the inter-node Round-Trip Time (RTT) is about 25 ± 0.1ms. For LAN experiments, the inter-node RTT is about 0.5 ± 0.03ms.
TABLE III: System Parameters.
       Parameters          Values
Number of sequencers 5, 10, 15, 20
Batch size for process transactions 500, 1000, 1500, 2000
The ratio of SGX, TDX, and CSV devices 1:2:2
We vary parameters in Table III to evaluate system gas cost, throughput, and latency. The experiment evaluates on-chain costs, as well as the throughput and latency of off-chain transaction processing. Specifically, gas cost measures the expense of executing TSC functions, throughput quantifies the number of transactions handled by sequencers per second, and latency reflects the time required for sequencers to vote and process transactions.
VII-B On-Chain Cost of Functions
The on-chain execution costs of TeeRollup are measured in gas, which is a unit that measures the computational effort required to execute operations on Ethereum [49]. The gas consumption depends on the transaction size and execution costs of smart contracts. To make these costs more understandable, we convert gas into USD with a gas price of 19.26 Gwei ($1.926\times 10^{-8}$ ETH) and an ETH price of 3376.77 USD on January 1, 2025 [53, 54].
TABLE IV: Protocol Execution Cost. Execution costs are estimated base on the prices (ETH/USD 3376.77 and gas price is 19.26 Gwei) on Jan. 1, 2025 [54, 53]
.      Methods Cost Gas USD Deposit 48551 3.16 UpdateState 156,263 10.16 StartChallenge 47,118 3.06 ResolveChallenge 146,618 9.53 SettleRollup 29,078 1.89 SettleWithdraw 124,511 8.10 Simple ETH transfer 21,000 1.37
Table IV shows our excellent performance in terms of on-chain cost. We measure the complete process of deposit, state update, and challenge on Ethereum when the number of sequencers is 10 (see Sec. V) and the batch size for process transactions is 2000. The UpdateState is one of the most gas-consuming methods, which costs about 156K gas (10.16 USD). That is because it includes the verification of the QC and the storage of new state on-chain. However, since the cost is amortized over 2000 transactions, the fee per transaction is reduced to just 78 gas (0.005 USD), which is significantly lower than the gas consumption of a simple ETH transfer (1.37 USD).
In terms of the challenge, a challenge-resolve process consists of the StartChallenge and ResolveChallenge, which totally require 194K gas (12.59 USD). The SettleRollup and SettleWithdraw, which are crucial for the redemption of the client, totally consuming 154K gas (9.99 USD). However, this only occurs when the challenge fails. In normal cases, the client’s withdrawal only requires a transfer in TeeRollup, which costs 78 gas (0.005 USD).
VII-C Performance Evaluation
We evaluate the performance of TeeRollup by measuring its throughput and latency under varying numbers of sequencers and batch sizes. We also evaluate the impact of introduced TEE by comparing throughput and latency with sequencers enabled and disabled with TEE.
Throughput and latency. Fig. 3 illustrates the throughput and latency of TeeRollup across varying numbers of sequencers and batch sizes in both LAN and WAN environments. In the LAN environment, as the number of sequencers increases, throughput experiences a slight decrease, while latency rises. This behavior is primarily due to the increased messaging overhead among sequencers as their number grows. Throughput improves with larger batch sizes, as they enable TeeRollup to process more transactions concurrently. For instance, with 5 sequencers and a batch size of 2000, throughput can reach about 28 KTPS. However, larger batch sizes also result in higher latency, which peaks at 119 ms. This increase in latency occurs because sequencers need more time to process the larger volume of transactions, leading to longer delays.
In the WAN environment, throughput and latency exhibit similar trends to those in the LAN environment. However, due to the higher communication delays inherent in WAN, both throughput is lower and latency is higher compared to the LAN setting. Furthermore, as batch size increases, the rise in latency in the WAN environment is less pronounced than in the LAN environment. This difference occurs because, in the WAN, the primary factor influencing latency shifts from transaction processing delays to communication delays.
(a) Throughput in LAN
(b) Latency in LAN
(c) Throughput in WAN
(d) Latency in WAN
Figure 3: Transaction processing performance with varying different numbers of sequencers and batch size in LAN and WAN.
TEE overhead. Fig. 4 show the influence of introducing TEE on throughput and latency in LAN and WAN, respectively. The operations executed inside a TEE need encryption and to switch context from the regular execution environment of the host machine, which degrades performance. Thus, the protection provided by TEEs reduces throughput and increases latency. For example, in the WAN environment with 20 sequencers, throughput decreases by 15.74%, while latency increases by 20.17%. Importantly, even with TEEs’ hardware protection in WAN, the reduction in performance does not exceed the acceptable range for users, maintaining a fine balance between security and performance.
(a) Throughput in LAN
(b) Latency in LAN
(c) Throughput in WAN
(d) Latency in WAN
Figure 4: System performance with/without SEV in LAN and WAN.
VII-D Comparison with Counterparts
We compare TeeRollup with existing rollup schemes, including StarkNet [10], Scroll [11], Optimism [4], and Arbitrum [7]. The average transaction fees for StarkNet and Scroll over the past month were approximately 0.043 USD and 0.114 USD, respectively [6]. The average transaction fees for Optimism and Arbitrum were significantly lower, at 0.012 USD and 0.005 USD, respectively. As shown in Sec. VII-B, the average transaction fee for TeeRollup is 0.006 USD per transaction with a batch size of 2000 (matching that of StarkNet), representing an 86% reduction compared to StarkNet. Thus, the transaction fee in TeeRollup is comparable to those of optimistic rollups. As for throughput, current rollup solutions are constrained by the Ethereum mainnet, which can support up to 3,000 TPS. As shown in Sec. VII-C, TeeRollup can achieve throughput exceeding 5000 TPS. Regarding withdrawal times, StarkNet and Scroll offer withdrawals within a few minutes, while Optimism and Arbitrum have a withdrawal period of one week. Although TeeRollup must undergo a challenge process when TEEs are unavailable, the withdrawal time during the normal case remains a few minutes.
VIII Conclusion
In this paper, we introduce TeeRollup, a high-performance and cost-effective rollup solution that leverages TEEs to optimize the gas cost and withdraw delay of rollup. TeeRollup employs a group of sequencers, safeguarded by heterogeneous TEEs, to process transactions off-chain while leveraging DAPs to ensure data availability. TeeRollup also adopts a cost-efficient challenge mechanism to ensure redeemability, even in scenarios where all TEEs become unavailable. Additionally, TeeRollup features a laziness penalty mechanism to encourage DAPs to work diligently. The prototype implementation of TeeRollup demonstrates promising results, outperforming state-of-the-art approaches in both on-chain cost efficiency and off-chain processing speed.